### v4

Abstract
Many real objects are modeled as discrete sets of points, such as corners or other salient features. For our main applications in chemistry, points represent atomic centers in a molecule or a solid material. We study the problem of classifying discrete (finite and periodic) sets of unordered points under isometry, which is any transformation preserving distances in a metric space.
Experimental noise motivates the new practical requirement to make such invariants Lipschitz continuous so that perturbing every point in its $\varepsilon$-neighborhood changes the invariant up to a constant multiple of $\varepsilon$ in a suitable distance satisfying all metric axioms. Since the given points are unordered, the key challenge is to compute all invariants and metrics in a near-linear time of the input size.
We define the Pointwise Distance Distribution (PDD) for any discrete set and prove, in addition to the properties above, the completeness of PDD for all periodic sets in general position. The PDD can compare nearly 2 million crystals from the world’s five largest databases within 2 hours on a modest desktop computer. The impact is upholding data integrity in crystallography because the PDD will not allow anyone to claim a ‘new’ material as a noisy disguise of a known crystal.
keywords:
isometry classification, complete invariant, continuous metric, periodic crystal
{MSCcodes}
74E15, 68U05, 51N20
1 Introduction: motivations, problem statement, and contributions
This paper is a substantial extension of the 10-page conference version at NeurIPS 2022 [widdowson2022resolving]. The original paper introduced the Pointwise Distance Distribution (PDD) as an isometry invariant of a periodic set of points in any Euclidean space $\mathbb{R}^{n}$, and claimed the key properties (Lipschitz continuity, near-linear time computability, and generic completeness) in the fully periodic case, where proofs in the appendices of [widdowson2022resolving] were not expected to be reviewed. This extended version defines PDD for any discrete set in a metric space and rigorously proves the properties above in finite and $l$-periodic cases for all $l\leq n$. We adapt the invariants to a more convenient form, speed up the original implementation almost by two orders of magnitude, and report new experiments revealing duplicates in the world’s largest materials databases.
The continuous and generically complete invariants are motivated by the previously unresolved ambiguity of digital representations of molecules and crystals in terms of atomic coordinates or lattice bases. Fig. 1 (middle) shows that the same periodic set can be obtained by periodically repeating different motifs of points.
Figure 1: Left: a lattice can be defined by many primitive bases. Middle: a periodic set can be defined by different pairs (basis, motif). Right: a hierarchy of discrete sets, which model periodic crystals and amorphous solids with points at atomic centers, see Definitions 1.11.21.5, and 3.4.
The crucial question “same or different?” was explicitly raised for crystals [sacchi2020same] and makes sense for many other real objects. For a cloud of unordered points in computer vision or chemistry applications, a list of atomic coordinates depends on a given coordinate system and an order of atoms. The independence of coordinate representations is important for identifying rigid structures and rigid conformations of flexible molecules such as proteins whose properties depend on a rigid shape.
Noisy measurements imply that any real objects are at least slightly different. Hence the next practical question is “how much different?” If noise is ignored up to any positive threshold, noisy perturbations of atomic centers can be continued sufficiently long to make any given sets identical. This sorites paradox [hyde2011sorites] can be resolved by quantifying even tiny differences through a continuous distance metric.
Definition 1.1 (a discrete set $S$ in a metric space $X$ with a metric $d_{X}$).
A metric space is any set $X$ of objects (called points) with a distance metric $d:X\times X\to\mathbb{R}$ satisfying the metric axioms: (1) coincidence $d_{X}(a,b)=0$ if and only if $a=b$, (2) symmetry $d_{X}(a,b)=d_{X}(b,a)$, and (3) triangle inequality $d_{X}(a,b)+d_{X}(b,a)\geq d_{X}(a,c)$ for any points $a,b,c\in X$. A set $S\subset X$ is called discrete if there is a constant $\varepsilon>0$ such that all points of $S$ are $\varepsilon$-separated, so $d_{X}(a,b)\geq\varepsilon$ for any $a,b\in S$.
An example of a discrete set $S$ is a finite set in $\mathbb{R}^{n}$ with the Euclidean metric denoted by $|\vec{p}-\vec{q}|$ for any points $p,q\in\mathbb{R}^{n}$. Here $\vec{p}$ denotes the vector from the origin $0\in\mathbb{R}^{n}$ to $p$. The positivity $d_{X}(a,b)\geq 0$ follows from other axioms: $2d_{X}(a,b)=d_{X}(a,b)+d_{X}(b,a)\geq d_{X}(a,a)=0$. Without the first axiom, $d$ is called a pseudo-metric and can be the zero function: $d_{X}(a,b)=0$ for all $a,b$. If the triangle inequality is allowed to fail with any additive error $\varepsilon>0$, the results of clustering such as $k$-means and DBSCAN can be predetermined and hence may not be trustworthy [rass2024metricizing].
Definition 1.2 (lattice, unit cell, motif, $l$-periodic set).
Vectors $\vec{v}_{1},\dots,\vec{v}_{n}\in\mathbb{R}^{n}$ form a basis if any vector in $\mathbb{R}^{n}$ can be written as $\vec{v}=\sum\limits_{i=1}^{n}x_{i}\vec{v}_{i}$ for unique $x_{1},\dots,x_{n}\in\mathbb{R}$. For $1\leq l\leq n$, the vectors $\vec{v}_{1},\dots,\vec{v}_{l}$ define the lattice $\Lambda=\{\sum\limits_{i=1}^{l}c_{i}\vec{v}_{i}\mid c_{1},\dots,c_{l}\in\mathbb{Z}\}$ and the unit cell $U=\{\sum\limits_{i=1}^{n}x_{i}\vec{v}_{i}\mid x_{1},\dots,x_{l}\in[0,1),x_{l+1},\dots,x_{n}\in\mathbb{R}\}\subset\mathbb{R}^{n}$. If $l=n$, then $U$ is an $n$-dimensional parallelepiped. If $l<n$, then $U$ is an infinite slab over an $l$-dimensional parallelepiped on $\vec{v}_{1},\dots,\vec{v}_{l}$. For any finite set of points (called a motif) $M\subset U$, the sum $S=M+\Lambda=\{\vec{p}+\vec{v}\mid p\in M,v\in\Lambda\}$ is an $l$-periodic point set.
Any unit cell $U$ includes only a partial boundary: we exclude the points with any coefficient $t_{i}=1$, $i=1,\dots,l$, for convenience. Then $\mathbb{R}^{n}$ for $l=n$ is tiled by the shifted cells $\{U+\vec{v}\mid\vec{v}\in\Lambda\}$ without overlaps. Any lattice is an example of a periodic set with one point in a motif. Any periodic point set $S=M+\Lambda$ can be considered a finite union $\bigcup_{p\in M}(\vec{p}+\Lambda)$ of lattices whose origins are shifted to all $p\in M=S\cap U$.
If we double a unit cell in one direction, e.g. by taking the basis $2\vec{v}_{1},\vec{v}_{2},\dots,\vec{v}_{n}$, the doubled motif $M\cup(M+\vec{v}_{1})$ with the sublattice on the new basis defines the original periodic point set $S=M+\Lambda$. A basis and its cell $U$ of $S$ are called primitive if $S\cap U$ has the smallest size among all unit cells $U$ of $S$. Fig. 1 (left) shows a square lattice in $\mathbb{R}^{2}$, which (as any lattice) can be generated by infinitely many primitive bases. Even if we fix a basis, Fig. 1 (middle) shows that different motifs in the same primitive cell $U$ define equivalent periodic sets, which differ only by translation.
Finite and periodic point sets represent molecules and periodic crystals at the atomic scale by considering zero-sized points at all atomic centers. Chemical bonds can be modelled by straight-line edges between atomic centers. However, even the strongest covalent bonds within a molecule depend on various thresholds for distances and angles. So these bonds are not real sticks and only abstractly represent inter-atomic interactions, while atomic nuclei are real objects. We model all materials at the fundamental level of atoms, which will suffice for all real materials. Since any object can be defined in many different ways, Definition 1.3 formalizes an equivalence.
Definition 1.3 (equivalence relation).
An equivalence is a binary relation (denoted by $\sim$) on any kind of objects satisfying the following axioms: (1) reflexivity: any objects $S$ is equivalent to itself, so $S\sim S$; (2) symmetry: if $S\sim Q$, then $Q\sim S$; (3) transitivity: if $S\sim Q$ and $Q\sim T$, then $S\sim T$. Any object $S$ defines its equivalence class $S]=\{Q\mid Q\sim S\}$ as the full collection of all objects $Q$ equivalent to $S$.
The transitivity axiom justifies that all equivalence classes are disjoint: if $[S]$ and $[T]$ share a common object $Q$, then $[S]=[T]$. Any well-defined classification should be based on an equivalence, whose practical examples are considered below.
Definition 1.4 (isometry, rigid motion in $\mathbb{R}^{n}$).
In a metric space $X$, an isometry is any map $f:X\to X$ that preserves inter-point distances, i.e. $d(f(p),f(q))=d(p,q)$ for all $p,q\in X$. In $\mathbb{R}^{n}$, any isometry decomposes into translations, rotations, and reflections, which generate the Euclidean group $\mathrm{E}(n)$. If reflections are excluded, orientation-preserving isometries are also called rigid motions and form group $\mathrm{SE}(n)$.
Rigid motion (denoted by $\cong$) is the strongest equivalence for many objects in practice because translations and rotations of a molecule or solid material keep all their properties at least under the same ambient conditions such as temperature and pressure. The isometry (denoted by $\simeq$) is only slightly weaker by allowing reflections. Taking compositions with a uniform scaling in $\mathbb{R}^{n}$ or including (say) affine transformations gives weaker equivalences that define smaller spaces of classes.
This paper focuses on isometry as a more general equivalence defined in any metric space. Our main problem will be to continuously parametrize equivalence classes of (various kinds of) discrete sets under isometry. Delone sets were introduced by B. Delone [delone1976local] as $(r,R)$-systems in $\mathbb{R}^{n}$ and make sense in any metric space $X$. Let $\bar{B}(p;r)=\{q\in X\mid d(p,q)\leq r\}$ be the closed ball with a center $p\in X$ and a radius $r$.
Definition 1.5 (Delone sets and $m$-regular sets).
In a metric space $X$, a Delone set $S$ is any subset of $X$ satisfying the following conditions:
(a) packing: there is a radius $r>0$ such that the closed balls $\bar{B}(p;r)$ for all points $p\in S$ are disjoint or, equivalently, all distances between points of $S$ are at least $2r$;
(b) covering: there is a radius $R>0$ such that $\bar{B}(p;R)$ for all $p\in S$ cover $X$, i.e. $\bigcup\limits_{p\in S}\bar{B}(p;R)=X$, or, equivalently, $\bar{B}(p;R)$ for any $p\in X$ has at least one point of $S$.
A Delone set is called $m$-regular if $S$ splits into $m$ classes under the global isometry equivalence: $p\sim q$ if there is an isometry $f:X\to X$ such that $f(S)=S$, $f(p)=q$.
The packing condition implies that $S$ is a discrete set in $X$ by specifying a minimum inter-point distance $\varepsilon=2r$ and is well-motivated by the fact that real atoms strongly repel each other at very short distances [feynman2011lectures]. The covering condition says that $X$ has no unbounded ‘empty’ balls without any points of $S$ and is also motivated by the absence of infinite round pores in solid materials, liquids, and dense gases.
All $m$-regular sets for $m>1$ are also called multi-regular, while 1-regular sets are often called regular. Any lattice $\Lambda\subset\mathbb{R}^{n}$ is regular because the required isometry $f:\Lambda\to\Lambda$ mapping a point $p\in\Lambda$ to another $q\in\Lambda$ is the translation by the vector $\vec{q}-\vec{p}$. Similarly, any periodic point set $S$ is $m$-regular, where $m$ is upper bounded by the size of a motif $M$ of $S$. A honeycomb periodic set in $\mathbb{R}^{2}$ modeling graphene is regular, but not a lattice because there are two points in a primitive unit cell. The regularity means that $S$ looks the same when viewed from any point of $S$. Fig. 1 (middle) shows a 2-regular set whose points split into red and blue classes under the global isometry equivalence. [dolbilin1998multiregular, Theorem 1.3] proved that any multi-regular Delone set is periodic.
A finite set in $\mathbb{R}^{n}$ is not a Delone set but any finite subset of a finite metric space is Delone. The latter special case is indicated by cyan and magenta regions slightly touching each other in Fig. 1 (middle). All other inclusions are strict, not to scale.
The key tool in classifying under an equivalence is an invariant that is a function $I$ taking the same value on all equivalent objects. For a finite set $S\subset\mathbb{R}^{n}$, the number $m$ of points is an isometry invariant, but the geometric average $\dfrac{1}{m}\sum\limits_{p\in S}\vec{p}$ is not.
We state the mapping problem for any discrete sets under isometry, though the same conditions make sense for many other objects, e.g. graphs and polygonal meshes, and equivalences, e.g. rigid motions, affine or projective transformations in $\mathbb{R}^{n}$.
Problem 1.6 (geo-mapping problem for any discrete sets under isometry).
For a metric space $X$ with a metric $d_{X}$, find a map $I:\{$discrete sets of unordered points in $X\}\to$ a metric space with a metric $d$ satisfying the following conditions.
(a) Completeness: any sets $S\simeq Q$ are isometric in $X$ if and only if $I(S)=I(Q)$.
(b) Realizability: the image $\{I(S)\mid S\subset X\}$ is parametrized so that taking any value of $I$ from this image allows us to reconstruct $S\subset X$ uniquely under isometry of $X$.
(c) Lipschitz continuity: there is a constant $\lambda$ such that if $Q$ is obtained by perturbing each point of $S$ up to any $\varepsilon$ in the metric $d_{X}$, then $d(I(S),I(Q))\leq\lambda\varepsilon$.
(d) Computability: the invariant $I$, the metric $d$, and the reconstruction of $S\subset X$ from $I(S)$ can be computed in a time that depends polynomially on the input sizes.
For any finite set $S\subset X$, its input size is the number $m$ of points. For any periodic point set $S\subset\mathbb{R}^{n}$, its input size is the number $m$ of points in a motif $M$ from Definition 1.2 because a Crystallographic Information File (CIF) specifying a basis and atomic coordinates in this basis has a linear length $O(m)$ in the motif size $m$. Some infinite Delone sets can described in a finite form, e.g. certain aperiodic crystals [senechal1996quasicrystals] can be obtained as projections of periodic crystals in higher dimensions.
We leave these general cases for future work and focus on finite and periodic point sets, which already cover many applications where Problem 1.6 was widely open.
Figure 2: Left: the symmetry group and a reduced cell discontinuously change under tiny noise. Middle: the space of 3 points under isometry is parametrized by inter-point distances $0<a\leq b\leq c\leq a+b$. Right: energy landscapes of crystals show optimized structures as isolated peaks of height==-energy. To see beyond the ‘fog’, we need a map parametrized by invariants in Problem 1.6.
The completeness in (1.6a) implies that the invariant $I$ is a descriptor with no false negatives and no false positives for all discrete sets, and hence can be considered a DNA-style code that uniquely identifies any isometry class. The realizability in (1.6b) is even stronger and enables us to sample the space of realizable invariants and reconstruct the resulting set $S$, while a real DNA code is insufficient to grow a living organism. The Lipschitz continuity in (1.6c) is motivated by ever-present thermal vibrations and experimental noise. Fig. 2 (left) shows that almost any perturbation of points can arbitrarily scale up a primitive cell. This inherent discontinuity of traditional cell-based representations remained a practical loophole in crystallography at least since 1965 [lawton1965reduced] and allowed disguising known materials by a slight perturbation changing the space group and even the primitive cell volume, and also by replacing some chemical elements to avoid detection by chemical composition [anosova2024importance, section 6].
Fig. 2 (middle) shows a solution of Problem 1.6 for $m=3$ points saying that any triangle is determined under isometry by 3 ordered inter-point distances. Real or simulated crystals are local optima (mountain peaks) in Fig. 2 (right) on a continuous space of (isometry classes of) periodic point sets, whose ‘geography’ was unknown.
Contributions. We introduce the Pointwise Distance Distribution for any discrete set in a metric space. This generality is of broad interest to experts in computational geometry and applications to physical objects from molecules to solid or even liquid materials. The previously unpublished aspects are the asymptotic for $l$-periodic sets, rigorous proofs of the Lipschitz continuity (also for adjusted and normalized invariants), near-linear time computability, and generic completeness in the finite and periodic case. The linear-time algorithms and the hierarchical nature of PDD computations have become extremely important for big databases, especially in the last years when millions of artificial structures were claimed ‘new’ without checking for duplication with known crystals. The decisive advance is closing this discontinuity loophole in crystallography, which is demonstrated for the world’s largest databases.
2 Review of rigorous approaches to mapping spaces of discrete sets
This section reviews progress in solving Problem 1.6 for finite and periodic point sets by proof-based methods, not by experimental studies, which are reviewed in [widdowson2022resolving, widdowson2023recognizing]. Finite sets have two subcases: ordered points (easy) and unordered (much harder).
Ordered finite sets. Kendall’s shape theory [kendall2009shape] studies ordered points $p_{1},\dots,p_{m}\in\mathbb{R}^{n}$ whose complete isometry invariant is the matrix of distances [schoenberg1935remarks] or scalar products $\vec{p}_{i}\cdot\vec{p}_{j}$ [weyl1946classical, chapter 2.9], see a linear-time invariant in [anosova2025complete, wlodawer2025duplicate]. A brute-force extension to $m$ unordered points requires $m!$ permutations of points, ruled out by (1.6d).
Unordered finite sets (point clouds). Extending the case of $m=3$ points in Fig. 2 (middle), Boutin and Kemper proved in 2004 that the unordered distribution of distances between $m$ points uniquely determines a generic $m$-point cloud $C\subset\mathbb{R}^{n}$ under isometry [boutin2004reconstructing]. The genericity condition allows almost all clouds apart from a measure 0 subspace among all clouds. For any cloud $C$ of $m$ unordered points in a metric space $X$, the vector $\mathrm{SPD}(C)$ consists of $\frac{m(m-1)}{2}$ Sorted Pairwise Distances written in increasing order and computable in time $O(m^{2}\log m)$. The space of 4-point clouds in $\mathbb{R}^{2}$ has dimension 5 because 6 inter-point distances satisfy one polynomial equation saying that the tetrahedron on these points has volume 0. Fig. 3 shows a 4-parameter family of pairs of non-isometric clouds with the same vector $\mathrm{SPD}$.
Figure 3: Non-isometric clouds of 4 points with the same 6 pairwise distances. Left: the trapezoid $T$ has points $(\pm 1,1)$, $(\pm 2,0)$. The kite $K$ has points $(3,0)$, (1,0)(-1,0), $(0,\pm 1)$. Right: the infinite family of non-isometric clouds $C^{+}\not\simeq C^{-}$ sharing $p_{1},p_{2},p_{3}$ and depending on parameters $a,b,c,d>0$ [caelli1979generating].
Problem 1.6 expands the question ‘Can we hear the shape of a drum?’ [kac1966can] which has the negative answer in terms of 2D polygons that are indistinguishable by spectral invariants [gordon1992isospectral, gordon1992one, reuter2006laplace, cosmo2019isospectralization, marin2021spectral]. Problem 1.6 looks for stronger invariants that can completely ‘sense’ as in (1.6b), not only ‘hear’, the rigid shape of any cloud.
Computational geometry studied weaker versions of Problem 1.6 by developing canonical representations of point clouds [alt1988congruence, brass2004testing], which can be considered complete invariants, and also metrics between isometry classes of clouds. For example, any metric between fixed clouds extends to their isometry classes [chew1999geometric] by minimization over infinitely many transformations from the group $\mathrm{E}(n)$. This extension of the Hausdorff distance [hausdorff1919dimension] for $m$-point clouds in $\mathbb{R}^{2}$ has time $O(m^{5}\log m)$, see [goodrich1999approximate]. The Gromov-Wasserstein metrics [memoli2011gromov, memoli2021gromov] are defined for any metric-measure spaces also by minimizing over infinitely many correspondences between points, but cannot be approximated with a factor less than 3 in polynomial time unless P=NP, see Corollary 3.8 in [schmiedl2017computational] and polynomial algorithms for partial cases in [agarwal2018computing, majhi2024approximating].
Computing a metric between isometry classes of clouds is only a small part of Problem 1.6. Indeed, to efficiently navigate on Earth, in addition to distances between cities, we need a satellite-type view of the full planet and hence a realizable continuous invariant $I$, which can be used as the latitude and longitude coordinates.
Geometric Data Science[anosova2025geometric, anosova2026seeing] gradually solved simpler versions of Problem 1.6 since 2020 when the continuity was first stated for lattices [mosca2020voronoi] and then for general periodic sets [anosova2021introduction, edelsbrunner2021density]. The case of 2D lattices was finished in [kurlin2024mathematics] with a weaker Hölder continuity for a stronger relation under rigid motion, because the Lipschitz continuity is impossible for perturbations of a lattice basis. See continuous chiral distances and geographic-style maps in [bright2023continuous, bright2023geographic], and complete invariants of 3D lattices in [kurlin2022complete, bright2021complete].
The Pointwise Distance Distribution (PDD) solved Problem 1.6 for finite unordered sets with distinct distances in $\mathbb{R}^{n}$ [widdowson2022resolving, Theorem 16]. This PDD appeared as a local distribution of distances in the finite case [memoli2011gromov] and was extended to higher order invariants in the periodic case [widdowson2025higher]. For all finite sets in $\mathbb{R}^{n}$, [widdowson2023recognizing, kurlin2024polynomial] developed complete invariants under rigid motion with Lipschitz continuous metrics. For periodic sets, the complete isoset [anosova2021isometry] has only an approximate algorithm for a continuous metric [anosova2026recognition, mcmanus2025computing]. For generic 1-periodic sets, [kurlin2025complete] defined an exact and continuous metric.
3 The Pointwise Distance Distribution and other isometry invariants
This section introduces the Pointwise Distance Distribution (PDD) for any discrete set $S$ with a specified finite subset $M$ in a metric space $X$. If $S$ is finite, we set $M=S$. If $S$ is periodic, $M$ is a motif of $S$, but $\mathrm{PDD}$ will depend only on $S$, not on $M$.
Definition 3.1 (PDD and AMD invariants).
Let $M=\{p_{1},\dots,p_{m}\}$ be a finite subset of a discrete set $S$ in a metric space $X$. Fix an integer $k\geq 1$. For every point $p\in M$, let $d_{1}(p)\leq\dots\leq d_{k}(p)$ be the distances from $p$ to its $k$ nearest neighbors within the full set $S$ (not restricted to $M$). The matrix $D(S,M;k)$ has $m$ rows consisting of the distances $d_{1}(p_{i}),\dots,d_{k}(p_{i})$ for $i=1,\dots,m$. If any $l\geq 2$ rows coincide, we collapse them into a single row with the weight $l/m$. The resulting unordered set (written as a matrix) of maximum $m$ rows and $k+1$ columns, including the extra column of weights, is the Pointwise Distance Distribution $\mathrm{PDD}(S,M;k)$. The Average Minimum Distance $\mathrm{AMD}_{i}$ is the weighted average of the $i$-th column in $\mathrm{PDD}(S,M;k)$ for each $i=1,\dots,k$. Let $\mathrm{AMD}(S,M;k)$ denote the vector $(\mathrm{AMD}_{1},\dots,\mathrm{AMD}_{k})$.
Definition 3.1 introduced the isometry invariant $\mathrm{PDD}(S,M;k)$ of a pair $(S,M)$ for a finite subset $M$ in any Delone set $S$. For any $l$-periodic point set $S\subset\mathbb{R}^{n}$, Theorem 3.3 will prove that $\mathrm{PDD}$ is independent of a motif $M\subset S$. We use the simpler notations $\mathrm{PDD}(S;k),\mathrm{AMD}(S;k)$ in the finite ($S=M$) and periodic cases.
Example 3.2 (4-point clouds $T,K$ in Fig. 3 (left)).
Table 1 shows the $4\times 3$ matrices $D(S;3)$ from Definition 3.1. The matrix $D(T;3)$ in Table 1 has two pairs of identical rows, so the matrix $\mathrm{PDD}(T;3)$ consists of two rows of weight $\frac{1}{2}$ below. The matrix $D(K;3)$ in Table 1 has only one pair of identical rows, so $\mathrm{PDD}(K;3)$ has three rows of weights $\frac{1}{2}$, $\frac{1}{4}$, $\frac{1}{4}$. Then $T,K$ are distinguished by $\mathrm{PDD}$s even for $k=1$.
Table 1: Each point of $T,K\subset\mathbb{R}^{2}$ in Figure 3 (left) has distances to other points in increasing order. After keeping only distances (not neighbors), the resulting $\mathrm{PDD}$s distinguish $T\not\simeq K$, see Example 3.2.
points of $T$ dist. to neighbor 1 dist. to neighbor 2 dist. to neighbor 3
(2,0)(-2,0) $\sqrt{2}$ to (1,+1)(-1,+1) $\sqrt{10}$ to $(+1,+1)$ 44 to $(+2,0)$
$(+2,0)$ $\sqrt{2}$ to $(+1,+1)$ $\sqrt{10}$ to (1,+1)(-1,+1) 44 to (2,0)(-2,0)
(1,+1)(-1,+1) $\sqrt{2}$ to (2,0)(-2,0) 22 to $(+1,+1)$ $\sqrt{10}$ to $(+2,0)$
$(+1,+1)$ $\sqrt{2}$ to $(+2,0)$ 22 to (1,+1)(-1,+1) $\sqrt{10}$ to (2,0)(-2,0)
points of $K$ dist. to neighbor 1 dist. to neighbor 2 dist. to neighbor 3
(1,0)(-1,0) $\sqrt{2}$ to (0,1)(0,-1) $\sqrt{2}$ to $(0,+1)$ 44 to $(3,0)$
$(+3,0)$ $\sqrt{10}$ to (0,1)(0,-1) $\sqrt{10}$ to $(0,+1)$ 44 to (1,0)(-1,0)
(0,1)(0,-1) $\sqrt{2}$ to (1,0)(-1,0) 22 to $(0,+1)$ $\sqrt{10}$ to $(3,0)$
$(0,+1)$ $\sqrt{2}$ to (1,0)(-1,0) 22 to (0,1)(0,-1) $\sqrt{10}$ to $(3,0)$
$\mathrm{PDD}(T)=\left(\begin{array}[]{c|ccc}1/2&\sqrt{2}&2&\sqrt{10}\\ 1/2&\sqrt{2}&\sqrt{10}&4\end{array}\right)\neq\mathrm{PDD}(K)=\left(\begin{array}[]{c|ccc}1/4&\sqrt{2}&\sqrt{2}&4\\ 1/2&\sqrt{2}&2&\sqrt{10}\\ 1/4&\sqrt{10}&\sqrt{10}&4\end{array}\right)$.
Theorem 3.3 extends [widdowson2022resolving, Theorem 3.2, proved in appendix C] from $n$-periodic sets to all finite sets, $l$-periodic sets, and arbitrary pairs $(S,M)$ from Definition 3.1.
Theorem 3.3 (invariance).
(a) For any isometry $S\to Q$ of discrete sets that maps a finite subset $M\subset S$ of $m$ points to $N\subset Q$, we have $\mathrm{PDD}(S,M;k)=\mathrm{PDD}(Q,N;k)$ and $\mathrm{AMD}(S,M;k)=\mathrm{AMD}(Q,N;k)$ for $1\leq k<m$. Hence, if $S=M$ is discrete, then $\mathrm{PDD}(S;k)$ and $\mathrm{AMD}(S;k)$ are isometry invariants of $S$.
(b) For any $l$-periodic point set $S\subset\mathbb{R}^{n}$, where $1\leq l\leq n$, $\mathrm{PDD}(S;k)$ and $\mathrm{AMD}(S;k)$ are invariants of $S$ (independent of a motif $M\subset S$) under isometry of $\mathbb{R}^{n}$ for $k\geq 1$.
Proof.
(a) For any sets $M\subset S$ and their isometric images $N\subset Q$, the invariance follows from the fact that any isometry preserves all inter-point distances.
(b) For any $l$-periodic point set $S=\Lambda+M\subset\mathbb{R}^{n}$, we first show that scaling up a cell $U$ and hence the motif $M=S\cap U$ of $m$ points keeps $\mathrm{PDD}$ invariant. For any integer $b\geq 1$, a matrix $B\in\mathrm{GL}(l;\mathbb{Z})$ with $|\det B|=b$ acts on the first $l$ vectors $\vec{v}_{1},\dots,\vec{v}_{l}$ that generate the $l$-dimensional base parallelepiped $P$ of $U$ in Definition 1.2.
Let $B(U)\subset\mathbb{R}^{n}$ denote the cell obtained from $U$ by applying $B$ to $P$ and keeping all other basis vectors $v_{l+1},\dots,v_{n}$ fixed. Then $D(S,S\cap B(U);k)$ from Definition 3.1 has the larger size $bm\times k$ but (due to periodicity of $S$) splits into $m$ blocks, each corresponding to $b$ points of the scaled motif $S\cap B(U)$ that are obtained from a single point $p\in M$ by translations by vectors of $\Lambda$. Since translations preserve distances, each of $m$ blocks has $b$ identical rows of distances to $k$ neighbors in $S$, the same as in $D(S,M;k)$. Then $\mathrm{PDD}(S,S\cap B(U);k)=\mathrm{PDD}(S,M;k)$ due to collapsing of identical rows in Definition 3.1. So $\mathrm{PDD}(S;k)$ is independent of any motif $M=S\cap U$.
Now we prove that $\mathrm{PDD}(S;k)$ is preserved by any isometry $f$ of $\mathbb{R}^{n}$. Any primitive cell $U$ of $S$ is bijectively mapped by $f$ to the unit cell $f(U)$ of $Q=f(S)$, which should be also primitive. Indeed, if $Q$ is preserved by a translation along a vector $v$ that doesn’t have all integer coefficients in the basis of $f(U)$, then $S=f^{-1}(Q)$ is preserved by the translation along $f^{-1}(v)$, which doesn’t have all integer coefficients in the basis of $U$, so $U$ was non-primitive. Since $U$ and $f(U)$ have the same number of points from $S$ and $Q=f(S)$, the isometry $f$ gives a bijection between the motifs of $S,Q$.
For any periodic sets $S,Q$, because $f$ maintains distances, every list of ordered distances from $p_{i}\in S\cap U$ to its first $k$ nearest neighbors in $S$ coincides with the list of the ordered distances from $f(p_{i})$ to its first $k$ neighbors in $Q$. These coincidences of distance lists give $\mathrm{PDD}(S;k)=\mathrm{PDD}(Q;k)$ after collapsing identical rows.
If we increase $k$, more columns with larger values are added to $\mathrm{PDD}(S;k)$ but all previous distances remain the same. Definition 3.4 will help describe the asymptotic of $\mathrm{PDD}(S;k)$ as $k\to+\infty$ in Theorem 3.7, which uses Lemma 3.5 extending [widdowson2022average, Lemma 11] to $l$-periodic sets $S\subset\mathbb{R}^{n}$ for any $1\leq l\leq n$, see all skipped proofs in appendix C
Definition 3.4 (Point Packing Coefficient $\mathrm{PPC}$ of a cell-periodic set $S$).
For $1\leq l\leq n$ and a basis $\vec{v}_{1},\dots,\vec{v}_{n}\in\mathbb{R}^{n}$, consider the lattice $\Lambda=\{\sum\limits_{i=1}^{l}c_{i}\vec{v}_{i}\mid c_{1},\dots,c_{l}\in\mathbb{Z}\}$ and the unit cell $U=\{\sum\limits_{i=1}^{n}x_{i}\vec{v}_{i}\mid x_{1},\dots,x_{l}\in[0,1),x_{l+1},\dots,x_{n}\in\mathbb{R}\}$. A discrete set $S\subset\mathbb{R}^{n}$ is cell-periodic if $S$ has a fixed number $m$ points in every shifted cell $U+\vec{v}$ for all $\vec{v}\in\Lambda$. If $l<n$, let $R^{l}\subset\mathbb{R}^{n}$ be the subspace spanned by $\vec{v}_{1},\dots,\vec{v}_{l}$, then $U$ is an infinite slab based on the $l$-dimensional parallelepiped of volume $\mathrm{vol}[U\cap R^{l}]$ . The volume of the unit ball in $\mathbb{R}^{l}$ is $V_{l}=\dfrac{\pi^{l/2}}{\Gamma(\frac{l}{2}+1)}$, where Euler’s Gamma function is $\Gamma(m)=(m-1)!$ and $\Gamma(\frac{m}{2}+1)=\sqrt{\pi}(m-\frac{1}{2})(m-\frac{3}{2})\cdots\frac{1}{2}$ for any integer $m\geq 1$. Define the Point Packing Coefficient of $S$ as $\mathrm{PPC}(S)=\sqrt[l]{\dfrac{\mathrm{vol}[U\cap R^{l}]}{mV_{l}}}$.
Any $l$-periodic set is cell-periodic, but all cell-periodic sets form a wider collection of Delone sets and model disordered solid materials that can have an underlying lattice with atoms at different positions in periodically translated cells $U+\vec{v}$, see Fig. 1.
Lemma 3.5 (bounds on points within a cylinder).
For any $1\leq l\leq n$ and a basis $\vec{v}_{1},\dots,\vec{v}_{n}\in\mathbb{R}^{n}$, let $S\subset\mathbb{R}^{n}$ be a cell-periodic set with a unit cell $U$ based on the $l$-dimensional parallelepiped $U\cap R^{l}$, where $R^{l}\subset\mathbb{R}^{n}$ is spanned by $\vec{v}_{1},\dots,\vec{v}_{l}$. Define the width $w$ of $U$ as $\sup\limits_{u,v\in U\cap R^{l}}|\vec{u}-\vec{v}|$. For any point $p\in S\cap U$ and a radius $r$, consider the cylinder $C(p;r)=\{\sum\limits_{i=1}^{n}x_{i}\vec{v}_{i}\text{ such that }x_{1},\dots,x_{n}\in\mathbb{R}\text{ and }|p-\sum\limits_{i=1}^{l}x_{i}\vec{v}_{i}|\leq r\}\subset\mathbb{R}^{n}$,
$\text{the \emph{lower} union }U^{-}(p;r)=\bigcup\{(U+\vec{v})\text{ such that }\vec{v}\in\Lambda,(U+\vec{v})\subset C(p;r)\}\subset\mathbb{R}^{n},$
$\text{the \emph{upper} union }U^{+}(p;r)=\bigcup\{(U+\vec{v})\text{ such that }\vec{v}\in\Lambda,(U+\vec{v})\cap C(p;r)\neq\emptyset\}.$
Let the unions $U^{\pm}(p;r)$ contain $m^{\pm}(p;r)$ shifted cells of $U+\vec{v}$ for some $\vec{v}\in\Lambda$. Let $S$ have $m=|S\cap U|$ points in $U$. Then the number of points from $S$ in $C(p;r)$ satisfies
$\left(\dfrac{r-w}{\mathrm{PPC}(S)}\right)^{l}\leq m^{-}(p;r)m\leq|S\cap C(p;r)|\leq m^{+}(p;r)m\leq\left(\dfrac{r+w}{\mathrm{PPC}(S)}\right)^{l}.$
Lemma 3.6 (distance bounds).
In the notations of Lemma 3.5, let the subspace $R^{n-l}$ be orthogonal to $R^{l}$, which is spanned by the first $l$ basis vectors of a cell $U$. Let the height $h$ of a cell-periodic set $S\subset\mathbb{R}^{n}$ with the cell $U$ be the maximum distance between points in the orthogonal projection of $S$ to $R^{n-l}$, so if $l=n$, then $h=0$. For any point $p\in S\cap U$, let $d_{k}(S;p)$ be the distance from $p$ to its $k$-th nearest neighbor in the full set $S$. Then $\mathrm{PPC}(S)\sqrt[l]{k}-w<d_{k}(S;p)\leq\sqrt{(\mathrm{PPC}(S)\sqrt[l]{k}+w)^{2}+h^{2}}$, $k\geq 1$.
Theorem 3.7 (asymptotic of $\mathrm{PDD}(S;k)$ as $k\to+\infty$).
For any point $p$ in a cell-periodic set $S\subset\mathbb{R}^{n}$, let $d_{k}(S;p)$ be the distance from $p$ to its $k$-th nearest neighbor in $S$. Then $\lim\limits_{k\to+\infty}\dfrac{d_{k}(S;p)}{\sqrt[l]{k}}=\mathrm{PPC}(S)$ and hence $\lim\limits_{k\to+\infty}\dfrac{\mathrm{AMD}_{k}(S)}{\sqrt[l]{k}}=\mathrm{PPC}(S)$.
Proof of Theorem 3.7.
Lemma 3.6 gives the following bounds for $\delta_{k}=\dfrac{d_{k}(S;p)}{\sqrt[l]{k}}-\mathrm{PPC}(S)$. The lower bound is $\delta_{k}>-u_{k}$, where $u_{k}=\dfrac{w}{\sqrt[l]{k}}\to 0$ as $k\to+\infty$ because $w$ is fixed. The upper bound is $\delta_{k}\leq\sqrt{(\mathrm{PPC}(S)+u_{k})^{2}+(h/\sqrt[l]{k})^{2}}-\mathrm{PPC}(S)\to 0$ as $k\to+\infty$, because $h$ is fixed. Hence $\delta_{k}=\dfrac{d_{k}(S;p)}{\sqrt[l]{k}}-\mathrm{PPC}(S)\to 0$ as $k\to+\infty$.
By Theorem 3.7, $\mathrm{AMD}_{k}(S)$ and all distances in the last column of $\mathrm{PDD}(S;k)$ asymptotically approach $\mathrm{PPC}(S)\sqrt[l]{k}$ as $k\to+\infty$ and hence are mainly determined by $\mathrm{PPC}(S)$ for large $k$. That is why the most descriptive information is contained in $\mathrm{PDD}(S;k)$ for smaller values of $k$, e.g. we use $k=100$ atomic neighbors in most experiments on crystals. To neutralize the asymptotic growth, we subtract and also normalize by the term $\mathrm{PPC}(S)\sqrt[l]{k}$ to get simpler invariants under uniform scaling.
[…]
Let $S\subset\mathbb{R}^{n}$ be any $l$-periodic set with an underlying lattice generated by $l$ vectors. The Average Deviation from Asymptotic is $\mathrm{ADA}_{k}(S)=\mathrm{AMD}_{k}(S)-\mathrm{PPC}(S)\sqrt[l]{k}$ for $k\geq 1$. The Pointwise Deviation from Asymptotic $\mathrm{PDA}(S;k)$ is obtained from the matrix $\mathrm{PDD}(S;k)$ by subtracting $\mathrm{PPC}(S)\sqrt[l]{j}$ from any distance in a row $i$ and a column $j$ for $i\geq 1\leq j\leq k$. The Average Normalized Deviation is and $\and_{k}(S)=\mathrm{ADA}_{k}(S)/(\mathrm{PPC}(S)\sqrt[l]{k})$, $k\geq 1$. The Pointwise Normalized Deviation $\mathrm{PND}(S;k)$ obtained from $\mathrm{PDA}(S;k)$ by dividing every element in a row $i$ and a column $j$ by $\mathrm{PPC}(S)\sqrt[l]{j}$ for $i\geq 1\leq j\leq k$.
Corollary 3.9 (invariance of and $\and,\mathrm{PND}$ under uniform scaling).
For any $l$-periodic set $S\subset\mathbb{R}^{n}$, and $\and_{k}(S)$ and $\mathrm{PND}(S;k)$ in Definition 3.8 are invariant under isometry and uniform scaling for any $k\geq 1$. Moreover, and $\and_{k}(S)\to 0$ as $k\to+\infty$.
Proof.
By Theorem 3.3, $\mathrm{PDD}(S;k)$ and hence all deviations in Definition 3.8 are invariant under isometry. Under uniform scaling $p\mapsto cp$ for a real constant $c\neq 0$, any inter-point distance and $\mathrm{PPC}(S)=\sqrt[l]{\dfrac{\mathrm{vol}[U\cap R^{l}]}{mV_{l}}}$ is multiplied by $c$ because $\mathrm{vol}[U\cap R^{l}]$ is scaled by the factor $c^{l}$. Hence and $\and_{k}(S)$ and $\mathrm{PND}(S;k)$ are invariant under both isometry and uniform scaling. To prove that and $\and_{k}(S)\to 0$ as $k\to+\infty$, use Theorem 3.7: and $\and_{k}(S)=\dfrac{\mathrm{ADA}_{k}(S)}{\mathrm{PPC}(S)\sqrt[l]{k}}=\dfrac{\mathrm{AMD}_{k}(S)}{\mathrm{PPC}(S)\sqrt[l]{k}}-1\to\dfrac{\mathrm{PPC}(S)}{\mathrm{PPC}(S)}-1=0$.
We conjecture that $\mathrm{ADA}_{k}(S)\to 0$ as $k\to+\infty$ without the extra division by $\sqrt[l]{k}$ for $l\geq 2$, which is confirmed by experiments on crystals and holds for $S=\mathbb{Z}^{n}$ by Example C.1.
The key input sizes for computing $\mathrm{PDD}(S;k)$ of any $l$-periodic point set $S\subset\mathbb{R}^{n}$ are the number $m$ of points in a unit cell $U$ and the number $k$ of neighbors. The full input consists of $k$, a basis of $U$ and a motif of $m$ points with coordinates in this basis as described in Definition 1.2. For a fixed dimension $n$ and other parameters, the asymptotic complexity of $\mathrm{PDD}(S;k)$ will depend near linearly on both $k,m$.
The output $\mathrm{PDD}(S;k)$ is a matrix with at most $m$ rows and exactly $k+1$ columns, where $m$ is the number of motif points. The first column contains the weights of rows, which sum to 1 and are proportional to the number of appearances of each row before collapsing in Definition 3.1, see Python code in appendix B. A different early version of Theorem 3.10 had a 5-line proof in [widdowson2022resolving, Theorem 5.1, appendix C].
Theorem 3.10 ($\mathrm{PDD}$ complexity).
Let $S\subset\mathbb{R}^{n}$ be any $l$-periodic set with a minimum inter-point distance $d_{\min}$ and a unit cell $U=P\times R^{n-l}$, where $P$ is a parallelepiped in the $l$-dimensional subspace $R^{l}$ with the orthogonal subspace $R^{n-l}$ in $\mathbb{R}^{n}$. Consider the width $w=\sup\limits_{u,v\in P}|\vec{u}-\vec{v}|$ and the height $h$ equal to the maximum distance between points in the orthogonal projection of $S$ to $R^{n-l}$. If the motif $M=S\cap U$ consists of $m$ points, then $\mathrm{PDD}(S;k)$ can be computed for any $k\geq 1$ in time
$O(km(2^{4n}\log k+\log m)+2^{12n}m\log^{2}k+(2^{8n}/l)k\log k+2^{8n}a^{l}bk),$
where $a=1+\dfrac{2.5w+2h}{\mathrm{PPC}(S)}$ and $b=\log(2\mathrm{PPC}(S)+3w+5h)-\log d_{\min}$. The complexity of $\mathrm{AMD}(S;k)$ and invariants $\mathrm{PDA}(S;k),\mathrm{PND}(S;k)$ from Definition 3.8 is the same as for $\mathrm{PDD}(S;k)$, because the extra computations can be done in time $O(km)$.
Proof of Theorem 3.10.
In the notations of Lemma 3.5, we have integers $1\leq l\leq n$ and a basis $\vec{v}_{1},\dots,\vec{v}_{n}$ of $\mathbb{R}^{n}$. The first $l$ basis vectors $\vec{v}_{1},\dots,\vec{v}_{l}$ generate the subspace $\mathbb{R}^{l}\subset\mathbb{R}^{n}$ and the lattice $\Lambda\subset\mathbb{R}^{l}$. Fix the origin $0\in\mathbb{R}^{n}$ at the center of the parallelepiped $U\cap R^{l}$. Then any point $p\in M=S\cap U$ is covered by the closed ball $\bar{B}(0;r)$ for the radius $r=\sqrt{(0.5w)^{2}+h^{2}}\leq 0.5w+h$. By Lemma 3.6, all $k$ neighbors of $p$ are covered by the closed cylinder $C(0;R)$ of the radius $R=r+\sqrt{(\mathrm{PPC}(S)\sqrt[l]{k}+w)^{2}+h^{2}}\leq\mathrm{PPC}(S)\sqrt[l]{k}+1.5w+2h$. To generate all $\Lambda$-translates of $M$ within $C(0;R)$, we gradually extend $U$ in cylindrical layers by adding more shifted cells $U+\vec{v}$ for vectors $v\in\Lambda$ until we get the upper union $U^{+}(0;R)$ covering the cylinder $C(0;R)$. The upper union $U^{+}(0;R)$ includes $k$ neighbors of each motif point and has the size $\nu=|S\cap U^{+}(0;R)|=m^{+}(0;R)m$ estimated by Lemma 3.5:
$\nu\leq\left(\dfrac{R+w}{\mathrm{PPC}(S)}\right)^{l}\leq\left(\dfrac{\mathrm{PPC}(S)\sqrt[l]{k}+2.5w+2h}{\mathrm{PPC}(S)}\right)^{l}=\left(\sqrt[l]{k}+\dfrac{2.5w+2h}{\mathrm{PPC}(S)}\right)^{l}=$
$=k\left(1+\dfrac{2.5w+2h}{\mathrm{PPC}(S)\sqrt[l]{k}}\right)^{l}\leq k\left(1+\dfrac{2.5w+2h}{\mathrm{PPC}(S)}\right)^{l}=a^{l}k,\text{ where }a=1+\dfrac{2.5w+2h}{\mathrm{PPC}(S)}.$
For the nearest neighbor search [elkin2022counterexamples], we can build a compressed cover tree on $\nu$ points of $T=S\cap U^{+}(0;R)$ in time $O(\nu c_{\min}^{8}\log\frac{2R+h}{d_{\min}})$ by [elkin2023new, Theorem 3.7], where $c_{\min}\leq 2^{n}$ is the minimized expansion constant of $T$, and $\frac{2R+h}{d_{\min}}$ is the upper bound for the ratio of max/min inter-point distances. Then $R\leq\mathrm{PPC}(S)\sqrt[l]{k}+1.5w+2h$ gives
$\log(2R+h)\leq\log(\sqrt[l]{k}(2\mathrm{PPC}(S)+3w+5h))=\log(2\mathrm{PPC}(S)+3w+5h)+(\log k)/l,$
$\text{so }\log\dfrac{2R+h}{d_{\min}}\leq b+\dfrac{1}{l}\log k,\text{ where }b=\log(2\mathrm{PPC}(S)+3w+5h)-\log d_{\min}.$
Using a compressed cover tree on $T$ by [elkin2023new, Theorem 4.9], we can find $k$ neighbors of $m$ points from $S\cap U$ among $\nu$ points of $T$ in time $O\big(mc^{2}(\log k)(c_{\min}^{10}\log\nu+ck)\big)$, where $c_{\min}\leq c\leq 2^{n}$ are expansion constants of $T$. Since $\log\nu\leq\log k+l\log a$, we compute distances from each of $m$ points to their $k$ nearest neighbors in $T$ in time
$O\big(\nu c_{\min}^{8}(b+(\log k)/l)\big)+O\big(mc^{2}\log k(c_{\min}^{10}\log\nu+ck)\big)\leq$
$O\big(a^{l}k2^{8n}(b+(\log k)/l)\big)+O\big(m2^{2n}\log k(2^{10n}(\log k+l\log a)+2^{2n}k)\big)\leq$
$O\big(2^{8n}a^{l}bk+(2^{8n}/l)k\log k\big)+O\big(2^{4n}m(k\log k+2^{8n}(\log^{2}k+l(\log a)\log k))\leq$
$O\big(2^{4n}(m+2^{4n}/l)k\log k+2^{12n}m\log^{2}k+2^{8n}a^{l}bk\big),\text{ we used }l\log a\leq O(\log k).$
The ordered lists of distances from points $p\in S\cap U$ to their $k$ nearest neighbors in $T$ are the rows of the matrix $D(S;k)$. If convenient, we can lexicographically sort $m$ lists of $k$ ordered distances, which needs time $O(km\log m)$, because a comparison of ordered lists of the length $k$ takes $O(k)$ time. The total time for $\mathrm{PDD}(S;k)$ is
$O(2^{4n}(m+2^{4n}/l)k\log k+2^{12n}m\log^{2}k+2^{8n}a^{l}bk)+O(km\log m)=$
$O(km(2^{4n}\log k+\log m)+2^{12n}m\log^{2}k+(2^{8n}/l)k\log k+2^{8n}a^{l}bk).$
The worst-case estimate in Theorem 3.10 is conservative due to the upper bound $2^{n}$ for the expansion constants $c_{\min},c$ from [elkin2023new, Definition 1.4]. We conjecture that this upper bound can be reduced to $2^{l}$ for any $l$-periodic point set $S\subset\mathbb{R}^{n}$.
For any fixed dimensions $l\leq n$, if we ignore the parameters $a,b,d_{\min}$, and $\mathrm{PPC}(S)$, then the complexity in Theorem 3.10 becomes $O(km(\log k+\log m))$, which is near-linear in both $k,m$. For the most practical dimensions $l=n=3$, experiments in section 6 will report running times in minutes on a modest desktop computer.
4 Lipschitz continuous Earth Mover’s Distance on invariants
[…]
Definition 4.1 of $\mathrm{EMD}$ makes sense for any matrix invariant $I(S)$ that is an unordered collection of row vectors $\vec{R}_{i}(S)$ with weights $w_{i}(S)\in(0,1]$ satisfying $\sum\limits_{i=1}^{m(S)}w_{i}(S)=1$. Each row $\vec{R}_{i}(S)$ should have a size independent of $i$, e.g. the number $k$ of neighbors in $\mathrm{PDD}(S;k)$. For any $\vec{R}_{i}=(r_{i1},\dots,r_{ik})$ and $\vec{R}_{j}=(r_{j1},\dots,r_{jk})$, the Minkowski distance is $L_{q}(\vec{R}_{i},\vec{R}_{j})=\big(\sum\limits_{l=1}^{k}|r_{il}-r_{jl}|^{q}\big)^{1/q}$, $L_{\infty}(\vec{R}_{i},\vec{R}_{j})=\max\limits_{l=1,\dots,k}|r_{il}-r_{jl}|$. We illustrate $\mathrm{EMD}$ for perturbations that scale up a unit cell, as in Fig. 2 (left). The integer sequence $\mathbb{Z}$ has $\mathrm{PDD}(\mathbb{Z};2)=(1;1,1)$, a single row of weight 1 and unit distances to 2 neighbors. The periodic sequence $\mathbb{Z}_{\varepsilon}=\{0,1+\varepsilon,2-\varepsilon\}+3\mathbb{Z}$ is obtained from $\mathbb{Z}$ by $\varepsilon$-perturbations of points $1,2$ and all their translates with period 3. Then $\mathrm{PDD}(\mathbb{Z}_{\varepsilon};2)=\left(\begin{array}[]{c|cccc}1/3&1+\varepsilon&1+\varepsilon\\ 2/3&1-2\varepsilon&1+\varepsilon\end{array}\right)$, where the 2nd row represents the shifted points $1+\varepsilon,2-\varepsilon$. After splitting $\mathrm{PDD}(\mathbb{Z};2)=(1;1,1)$ into two identical rows of weights $\frac{1}{3},\frac{2}{3}$ and using $L_{\infty}$ on vectors of two distances, a difference between $\mathrm{PDD}$s can be defined as the weighted average $\frac{1}{3}\varepsilon+\frac{2}{3}2\varepsilon=\frac{5}{3}\varepsilon$, which is generalized below.
Definition 4.1 (Earth Mover’s Distance $\mathrm{EMD}_{q}$).
Let discrete sets $S,Q$ in a metric space have weighted distributions $I(S),I(Q)$ as above. For any real $q\geq 1$, the Earth Mover’s Distance $\mathrm{EMD}_{q}(I(S),I(Q))=\sum\limits_{i=1}^{m(S)}\sum\limits_{j=1}^{m(Q)}f_{ij}L_{q}(\vec{R}_{i}(S),\vec{R}_{j}(Q))$ is minimized over variable parameters $f_{ij}\in[0,1]$ subject to the conditions $\sum\limits_{j=1}^{m(Q)}f_{ij}=w_{i}(S)$ for $i=1,\dots,m(S)$ and $\sum\limits_{i=1}^{m(S)}f_{ij}=w_{j}(Q)$ for $j=1,\dots,m(Q)$.
The first condition $\sum\limits_{j=1}^{m(Q)}f_{ij}\leq w_{i}(S)$ means that not more than the weight $w_{i}(S)$ of the vector $\vec{R}_{i}(S)$ ‘flows’ into all vectors $\vec{R}_{j}(Q)$ via optimized parameters $f_{ij}\in[0,1]$ for $j=1,\dots,m(Q)$. The second condition $\sum\limits_{i=1}^{m(S)}f_{ij}=w_{j}(Q)$ means that all ‘flows’ $f_{ij}$ from $\vec{R}_{i}(S)$ for $i=1,\dots,m(S)$ ‘flow’ into $\vec{R}_{j}(Q)$ up to the maximum weight $w_{j}(Q)$.
The EMD satisfies all metric axioms [rubner2000earth, appendix], needs $O(m^{3}\log m)$ time for distributions of a maximum size $m$, and can be approximated in $O(m)$ time [shirdhonkar2008approximate].
The Lipschitz continuity of invariants in $\mathrm{EMD}$ will use bounded perturbations of points up to $\varepsilon$ in the metric $d_{X}$ of an ambient space $X$. Since atoms are not outliers or noise, such perturbations can be formalized as the bottleneck distance $d_{B}(S,Q)=\inf\limits_{g:S\to Q}\;\sup\limits_{p\in S}d_{X}(g(p),p)$ minimized over all bijections $g:S\to Q$ between (possibly infinite) sets. This definition is computationally intractable even for finite sets due to exponentially many $m!$ bijections between sets of $m$ points. [widdowson2022resolving, Example 2.1] shows that the 1-dimensional lattices $\mathbb{Z}$ and $(1+\delta)\mathbb{Z}$ have $d_{B}=+\infty$ for any $\delta>0$.
If $S,Q$ are lattices of equal density (equal unit cell volume), they have a finite bottleneck distance $d_{B}$ by [duneau1991bounded, Theorem 1(iii)]. If we consider only periodic point sets $S,Q\subset\mathbb{R}^{n}$ with the same density (or unit cells of the same volume), $d_{B}(S,Q)$ becomes a well-defined wobbling distance [carstens1999geometrical], which is still discontinuous under perturbations by [widdowson2022resolving, Example 2.2], see related results for non-periodic sets in [laczkovich1992uniformly].
Recall that the packing radius $r(S)$ is the minimum half-distance between any points of $S$, or $r(S)$ is the maximum radius $r$ to have disjoint open balls of radius $r$ centered at all points of $S$. Theorem 4.2 extends [widdowson2022resolving, Theorem 4.3, proved in appendix C] to finite and $l$-periodic sets, and distances based on any Minkowski metric $L_{q}$.
Theorem 4.2 (Lipschitz continuity).
Let $M$ be a finite subset of a discrete set $S$ in a space $X$ with a metric $d_{X}$. Let $Q$ and its finite subset $T$ be obtained from $S$ and $M$, respectively, by perturbing every point of $S$ up to $\varepsilon$ in the metric $d_{X}$. Fix any real $q\in[1,+\infty]$ and an integer $k\geq 1$. Interpret $\sqrt[q]{k}$ as 1 in the limit case $q=+\infty$.
(a) Then $\mathrm{EMD}_{q}(\mathrm{PDD}(S,M;k),\mathrm{PDD}(Q,T;k))\leq 2\varepsilon\sqrt[q]{k}$.
(b) If $S,Q$ are $l$-periodic and $\min\{r(S),r(Q)\}>\varepsilon$, then $\mathrm{PPC}(S)=\mathrm{PPC}(Q)$, and
$\mathrm{EMD}_{q}(\mathrm{PDA}(S;k),\mathrm{PDA}(Q;k))\leq 2\varepsilon\sqrt[q]{k}$, $\mathrm{EMD}_{q}(\mathrm{PND}(S;k),\mathrm{PND}(Q;k))\leq\dfrac{2\varepsilon\sqrt[q]{k}}{\mathrm{PPC}(S)}$.
Theorem 4.2 is proved in appendix C similar to [widdowson2022average, Lemma 8] for $q=+\infty$. All columns of $\mathrm{PDD},\mathrm{PDA},\mathrm{PND}$ are ordered by the index $k$ of neighbors. Though their rows are unordered (as points of a motif $M$), all such matrices even with different numbers of rows can be compared by Earth Mover’s Distance, or by any other metrics on weighted distributions, see Definition 4.1. We can simplify any $\mathrm{PDD}$ into a fixed-size matrix, which can be flattened into a vector, while keeping the continuity and almost all invariant data. Any distribution of $m$ unordered values can be reconstructed from its $m$ moments below. When all weights $w_{i}$ are rational as in our case, the distribution can be expanded to equal-weighted values $a_{1},\dots,a_{m}$. The $m$ moments can recover all $a_{1},\dots,a_{m}$ as roots of a degree $m$ polynomial whose coefficients are expressed via the $m$ moments [macdonald1998symmetric], e.g. any $a,b\in\mathbb{R}$ can be found from $a+b,a^{2}+b^{2}$ as the roots of $x^{2}-(a+b)x+ab$, where $ab=\frac{1}{2}((a+b)^{2}-(a^{2}+b^{2}))$.
Let $A$ be any unordered set of real numbers $a_{1},\dots,a_{m}$ with weights $w_{1},\dots,w_{m}$, respectively, such that $\sum\limits_{i=1}^{m}w_{i}=1$. For any integer $t\geq 1$, the $t$-th moment [keeping1995introduction, section 2.7] is $\mu_{t}(A)=\sqrt[t]{m^{1-t}\sum\limits_{i=1}^{m}w_{i}a_{i}^{t}}$, so $\mu_{1}(A)=\sum\limits_{i=1}^{m}w_{i}a_{i}$ is the usual average.
For any integer $t\geq 2$, we avoid subtracting $\mu_{1}$ from the numbers $a_{1},\dots,a_{m}$, which would convert $\mu_{2}$ into the standard deviation $\sigma$, and normalize by the factor $m^{(1/t)-1}$ to guarantee the continuity of moments with the Lipschitz constant $\lambda=2$.
Definition 4.3 ($t$-moments matrix $\mu^{(t)}$}).
Fix any integer $t\geq 1$. Let $I(S)$ be a matrix invariant of a cell-periodic set $S$. For every column $A$ of $I(S)$, consisting of unordered numbers with weights, write the column $(\mu_{1}(A),\dots,\mu_{t}(A))$. All new columns form the $t$-moments matrix $\mu^{(t)}[I(S)]$, which has $t$ canonically ordered rows.
For $t=1$, the $1\times k$ matrix $\mu^{(1)}[\mathrm{PDD}(S;k)]$ appeared in Definition 3.1 as the vector $\mathrm{AMD}(S;k)=(\mathrm{AMD}_{1},\dots,\mathrm{AMD}_{k})$. All rows and columns of the matrix $\mu^{(t)}[I(S)]$ are ordered, but this matrix is a bit weaker than $I(S)$ because each column can be reconstructed from its moments (for a large enough $t$) only up to permutation. However, we can flatten any matrix $\mu^{(t)}[I(S)]$ to a vector for machine learning [balasingham2024material, balasingham2024accelerating].
Theorem 4.4 extends [widdowson2022resolving, Theorem 4.2, proved in appendix C] to the new invariants $\mathrm{PDA},\mathrm{PND}$ of any finite and $l$-periodic sets for a Minkowski metric $L_{q}$, $q\geq 1$.
Theorem 4.4 (lower bounds of $\mathrm{EMD}$).
For finite or $l$-periodic sets $S,Q\subset\mathbb{R}^{n}$,
(a) $\mathrm{EMD}_{q}(\mathrm{PDD}(S;k),\mathrm{PDD}(Q;k))\geq L_{q}(\mathrm{AMD}(S;k),\mathrm{AMD}(Q;k))$;
(b) $\mathrm{EMD}_{q}(\mathrm{PDA}(S;k),\mathrm{PDA}(Q;k))\geq L_{q}(\mathrm{ADA}(S;k),\mathrm{ADA}(Q;k))$;
[…]
5 Generic completeness of Pointwise Distance Distributions
We prove the generic completeness in both finite (easy) and periodic (much harder) cases.
Theorem 5.1.
Any cloud $C\subset\mathbb{R}^{n}$ of $m$ unordered points with distinct inter-point distances can be reconstructed from $\mathrm{PDD}(C;m-1)$, uniquely under isometry.
Proof of Theorem 5.1.
Since all inter-point distances are distinct, every such distance |pq||p-q| between points $p,q\in C$ appears twice in $\mathrm{PDD}(C;m-1)$: once in the row of $p$ and once in the row of $q$. After choosing an arbitrary order of points, $\mathrm{PDD}(C;m-1)$ suffices to reconstruct the classical distance matrix on ordered points. This distance matrix determines $C\subset\mathbb{R}^{n}$ uniquely under isometry [kruskal1978multidimensional].
Conjecture 5.2 (completeness of $\mathrm{PDD}$ in $\mathbb{R}^{2}$).
Any cloud $C\subset\mathbb{R}^{2}$ of $m$ unordered points can be reconstructed from $\mathrm{PDD}(C;m-1)$ uniquely under isometry.
Theorem 5.3 (completeness of $\mathrm{PDD}$ for $m\leq 4$ points).
$\mathrm{PDD}(C;m-1)$ is a complete isometry invariant of all clouds $C\subset\mathbb{R}^{n}$ for any $m\leq 4$ unordered points.
For a periodic point set $S\subset\mathbb{R}^{n}$, the generic completeness of $\mathrm{PDD}$ is much harder because infinitely many distances between points of $S$ are repeated due to periodicity. We introduce a few auxiliary concepts to define distance-generic periodic sets later.
For any point $p$ in a lattice $\Lambda\subset\mathbb{R}^{n}$, the open Voronoi domain $V(\Lambda;p)=\{q\in\mathbb{R}^{n}\text{ such that }|q-p|<|q-p^{\prime}|\mbox{ for any }p^{\prime}\in\Lambda-p\}$ is the neighborhood of all points $q\in\mathbb{R}^{n}$ that are strictly closer to $p$ than to all other points $p^{\prime}$ of the lattice $\Lambda$ [smith2022practical].
The Voronoi domains $V(\Lambda;p)$ of different points $p\in\Lambda$ are disjoint translation copies of each other and their closures tile $\mathbb{R}^{n}$, so $\cup_{p\in\Lambda}\bar{V}(\Lambda;p)=\mathbb{R}^{n}$. For example, for a generic lattice $\Lambda\subset\mathbb{R}^{2}$, the domain $V(\Lambda;p)$ is a centrally symmetric hexagon.
Points $p,p^{\prime}\in\Lambda$ are Voronoi neighbors if their Voronoi domains share a boundary point, so $\bar{V}(\Lambda;p)\cap\bar{V}(\Lambda,p^{\prime})\neq\emptyset$. Below we always assume that any lattice $\Lambda$ is shifted to contain the origin 0, also any periodic point set $S=\Lambda+M$ has a point at 0.
Definition 5.4 (neighbor set $N(\Lambda)$ and basis distances).
For any lattice $\Lambda\subset\mathbb{R}^{n}$, the neighbor set of the origin 0 is $N(\Lambda)=\Lambda\cap\bar{B}(0;r)\setminus\{0\}$ for a minimum radius $r$ such that $N(\Lambda)$ is not contained in any affine (n1)(n-1)-dimensional subspace of $\mathbb{R}^{n}$, and $N(\Lambda)$ includes all $n+1$ nearest neighbors (within $\Lambda$) of any point $q\in V(\Lambda;0)$.
Consider all sets of unordered points $p_{1},\dots,p_{n}\in N(\Lambda)$ that are linearly independent, i.e. the vectors $\vec{p}_{1},\dots,\vec{p}_{n}$ form a linear basis of $\mathbb{R}^{n}$. For any point $q\in V(\Lambda;0)$, a lexicographically smallest list of distances $d_{1}(q)\leq\dots\leq d_{n}(q)$ from $q$ to a set of linearly independent points $p_{1},\dots,p_{n}\in N(\Lambda)$ is called the list of basis distances of $q$.
The linear independence of vectors $\vec{p}_{1},\dots,\vec{p}_{n}$ in Definition 5.4 guarantees that any point $q$ is uniquely determined in $\mathbb{R}^{n}$ by its distances $|q|,d_{1}(q),\dots,d_{n}(q)$ to $n+1$ neighbors $0,p_{1},\dots,p_{n}$, which are not in the same (n1)(n-1)-dimensional subspace.
Let $\Lambda$ be generated by $(2,0),(0,1)$. The Voronoi domain $V(\Lambda;0)$ is the rectangle $(-1,1)\times(-0.5,0,5)$. The neighbor set $N(\Lambda)\subset\Lambda$ includes the 3rd neighbors $(0,\pm 2)$ of the points $(0,\pm 0.4)\in V(\Lambda;0)$. Indeed, if in Definition 5.4 $\Lambda$ has a radius $r<2$, then $\Lambda\cap\bar{B}(0;r)\setminus\{0\}=\{(0,\pm 1)\}$ is in the 1-dimensional subspace ($y$-axis) of $\mathbb{R}^{2}$. For $q=(0,0.4)$, considering all pairs $(\vec{p}_{1},\vec{p}_{2})$ that generate $\mathbb{R}^{2}$ among the four possibilities $((0,\pm 1),(\pm 2,0))$, we find the basis distances $d_{1}(q)=0.6<d_{2}(q)=\sqrt{0.4^{2}+2^{2}}\approx 2.04$ for the 2nd and 3rd lattice neighbors $p_{1}=(0,1)$ and $p_{2}=(\pm 2,0)$ of $q$, respectively.
Lemma 5.5.
The neighbor set $N(\Lambda)$ of any lattice $\Lambda$ is covered by $\bar{B}(0;2R(\Lambda))$, where the covering radius $R(\Lambda)$ is the minimum $R>0$ such that $\cup_{p\in\Lambda}\bar{B}(p;R)=\mathbb{R}^{n}$.
Proof of Lemma 5.5.
Any point $p$ in the closure $\bar{V}(\Lambda;0)$ of the Voronoi domain has $n+1$ lattice neighbors (within $\Lambda$) among them the origin $0\in\Lambda$ and at least $2(2^{n}-1)$ Voronoi neighbors of 0 [conway1992low]. In $\mathbb{R}^{n}$, any vertex of the boundary of $V(\Lambda;0)$ is equidistant to at least $n+1$ points of $\Lambda$ (the origin 0 and its $n$ Voronoi neighbors). The longest of these distances to Voronoi neighbors is the covering radius $R(\Lambda)$. The ball $\bar{B}(0;2R(\Lambda))$ covers all Voronoi neighbors of 0 and hence the neighbor set $N(\Lambda)$.
Definition 5.6 (a distance-generic set).
A periodic point set $S=M+\Lambda\subset\mathbb{R}^{n}$ with the origin $0\in\Lambda\subset S$ is called distance-generic if the following conditions hold.
(5.6a) For any points $p,q\in S\cap V(\Lambda;0)$, the vectors $\vec{p},\vec{q}$ are not orthogonal.
(5.6b) For vectors $\vec{u},\vec{v}$ between any two pairs of points in $S$, if $|\vec{u}|=l|\vec{v}|\leq 2R(\Lambda)$ for $l=1,2$, then $\vec{u}=\pm l\vec{v}$ and $\vec{v}\in\Lambda$.
(5.6c) For any point $q\in S\cap V(\Lambda;0)$, let $d_{0}=|q|$ be its distance to the closest neighbor $p_{0}=0$ in $\Lambda$. Take any linearly independent points $p_{1},\dots,p_{n}\in N(\Lambda)$ and any distances $d_{1}\leq\dots\leq d_{n}$ from $q$ to some points in $S\cap\bar{B}(0;2R(\Lambda))$. The $n+1$ spheres $\partial B(p_{i};d_{i})$ can meet at a single point of $S\cap V(\Lambda;0)$ only if $d_{1}\leq\dots\leq d_{n}$ are the basis distances of $q$ and only for two tuples $p_{1},\dots,p_{n}\in N(\Lambda)$ related by $\vec{v}\mapsto-\vec{v}$.
Condition (5.6b) means that all inter-point distances are distinct apart from necessary exceptions due to periodicity. Since any periodic set $S=M+\Lambda\subset\mathbb{R}^{n}$ is invariant under translations along all vectors of $\Lambda$, condition (5.6b) for $|\vec{v}|\leq 2R(\Lambda)$ can be checked only for vectors from all points of $S$ in the original Voronoi domain $V(\Lambda;0)$ to all points in the domain $3V(\Lambda;0)$ extended by factor 3. Condition (5.6b) implies that $S$ has no points on the boundary $\partial V(\Lambda;0)$, because any such point is equidistant to points $0,v\in\Lambda$ and hence should belong to $\Lambda$. Let a lattice distance be the Euclidean distance from any $p\in M=S\cap V(\Lambda;0)$ to its lattice translate $p+\vec{v}$ for all $\vec{v}\in\Lambda$. Condition (5.6a) guarantees that only a lattice distance $d$ appears together with $2d$ (and possibly with higher multiples) in a row of $\mathrm{PDD}(S;k)$. Any such $d$ and its multiples are repeated twice in every row, because $\Lambda$ is centrally symmetric.
Lemma 5.7 (almost any periodic set is distance-generic).
Let $S=M+\Lambda\subset\mathbb{R}^{n}$ be any periodic point set. For any $\varepsilon>0$, one can perturb coordinates of a basis of $\Lambda$ and of points from $M$ up to $\varepsilon$ such that the resulting perturbation $S^{\prime}$ of $S$ is a distance-generic periodic point set in the sense of Definition 5.6.
Proof.
We can assume that the motif $M$ of $S$ is a subset of the open Voronoi domain $V(\Lambda;0)$ and include the origin 0. We show below that conditions (5.6a,b) define a codimension 1 discriminant (singular subspace) in the space of all parameters $P$ that are coordinates of points of $M$ and of basis vectors of $\Lambda$. In condition (5.6a), for any points $p,q\in V(\Lambda;0)$, the orthogonality is expressed as $f_{a}(p,q)=\vec{p}\cdot\vec{q}=\sum\limits_{i=1}^{n}p_{i}q_{i}=0$. In condition (5.6b), for any vectors $\vec{u},\vec{v}$ that join points of $S$, have a maximum length $2R(\Lambda)$, and satisfy $u\neq\pm l\vec{v}$ for $l=1,2$, the equality $|\vec{u}|=l|\vec{v}|$ can be written as $f_{b}(u,v)=\sum\limits_{i=1}^{n}u_{i}^{2}-l^{2}\sum\limits_{i=1}^{n}v_{i}^{2}=0$. So condition (5.6a) forbids a codimension 1 subspace defined by finitely many equations $f_{b}(u,v)=0$ for all $u,v$ above.
Similarly, condition (5.6c) can be written via polynomial equations in point coordinates. For any fixed radii $d_{0},\dots,d_{n}$, almost all $n+1$ spheres in $\mathbb{R}^{n}$, whose centers are not in any (n1)(n-1)-dimensional affine subspace, have no common points. Hence condition (5.6c) also forbids a codimension 1 subspace. All involved functions in equations above are continuous in the coordinates of points and basis vectors. Then a motif $M=S\cap V(\Lambda;0)$ and a basis of $\Lambda$ can be slightly perturbed to move $S$ to $S^{\prime}$ outside the union of all finitely many codimension 1 subspaces above. Hence any periodic point set $S$ can be made distance-generic by a small enough perturbation.
The size $m$ of a motif $M$ is an isometry invariant because any isometry maps $N$ to another hose motif of the same size. In dimensions $n=2,3$, any lattice $\Lambda$ can be reconstructed from its complete isometry invariants [kurlin2024mathematics, kurlin2022complete]. Theorem 5.8 reconstructs a periodic point set $S=M+\Lambda\subset\mathbb{R}^{n}$ in any dimension $n\geq 2$ from the invariant $I(S)$ consisting of $m$, $\mathrm{PDD}(S;k)$, and (complete invariants of) a lattice $\Lambda$ to satisfy completeness (1.6a) for distance-generic periodic sets $S\subset\mathbb{R}^{n}$. New Lemma 5.7 and the arguments below clarify the early unreviewed proof in [widdowson2022resolving, appendix C].
Theorem 5.8 (generic completeness of $\mathrm{PDD}$).
Let $S=M+\Lambda\subset\mathbb{R}^{n}$ be any distance-generic periodic set whose motif $M$ has $m$ points. Let $R(\Lambda)$ be the smallest radius $R$ such that all closed balls with centers $p\in\Lambda$ and radius $R$ cover $\mathbb{R}^{n}$. For any $k$ such that all distances in the last column of $\mathrm{PDD}(S;k)$ are larger than $2R(\Lambda)$, $S$ can be reconstructed from $m$, $\Lambda$, and $\mathrm{PDD}(S;k)$, uniquely under isometry in $\mathbb{R}^{n}$.
Proof.
The given number $m$ of points in a unit cell $U$ of $\Lambda$ is a common multiple of all denominators in rational weights of the rows in the given matrix $\mathrm{PDD}(S;k)$. Enlarge $\mathrm{PDD}(S;k)$ by replacing every row of a weight $w$ with the integer number $mw$ of identical rows having the same weight $\frac{1}{m}$. One can assume that the origin $0\in\Lambda$ belongs to the motif $M$ of $S$ and is represented by the first row of $\mathrm{PDD}(S;k)$.
If $\mathrm{PDD}(S;k)$ has $m\geq 2$ rows, we will reconstruct all other m1m-1 points of the periodic point set $S$ within the open Voronoi domain $V(\Lambda;0)$. No points of $S$ can be on the boundary of $V(\Lambda;0)$ due to condition (5.6b) on distinct distances.
Remove from each row of $\mathrm{PDD}(S;k)$ all lattice distances between any points of $\Lambda$. Then every remaining distance is between only points $p,q\in S$ such that $\vec{p}-\vec{q}\not\in\Lambda$. Take a unique point $q\in S\cap V(\Lambda;0)\setminus\{0\}$ that has the smallest distance $d_{0}=|q|$ to the origin and hence uniquely determined in the row of $q$ in $\mathrm{PDD}(S;k)$. Then we will look for $n$ basis distances $d_{1}<\dots<d_{n}$ from $q$ to its further $n$ lattice neighbors $p_{1},\dots,p_{n}\in N(\Lambda)\subset\Lambda-0$ such that $\vec{p}_{1},\dots,\vec{p}_{n}$ form a linear basis of $\mathbb{R}^{n}$. All basis distances $d_{0},\dots,d_{n}$ are distinct due to (5.6b). By Lemma 5.5 they appear once in both rows of the points $0,q\in S$ in $\mathrm{PDD}(S;k)$ after the shortest distance $d_{0}=|q|$.
Though the basis distances of $q$ may not be the $n$ smallest values appearing after $d_{0}=|q|$ in the first and second rows of $\mathrm{PDD}(S;k)$, we will try all subsequences $d_{1}<\dots<d_{n}$ of distinct distances shared by both rows. Similarly, we cannot be sure that $n$ closest neighbors of $q$ in $S\setminus\{0\}$ define linearly independent vectors of $\Lambda$.
Hence we try all linearly independent points $p_{1},\dots,p_{n}\in N(\Lambda)$. For all finitely many choices, we check if the $n+1$ spheres $\partial B(p_{i};d_{i})$ meet at a single point in $V(\Lambda;0)$, which will be the required point $q$. These (n1)(n-1)-dimensional spheres are 1D circles for $n=2$ and 2D spheres for $n=3$. Condition (5.6c) will guarantee below a reconstruction of $q$ as a single intersection of these $n+1$ spheres of dimension n1n-1.
The basis distances $d_{1}<\dots<d_{n}$ of $q$ should form the lexicographically smallest list among all lists of distances from $q$ to points $p_{1},\dots,p_{n}\in N(\Lambda)$. This smallest list emerges for at most two tuples of linearly independent points $p_{1},\dots,p_{n}\in N(\Lambda)$ related by the isometry $\vec{v}\mapsto-\vec{v}$, which preserves $\Lambda$. For a first reconstruction outside $\Lambda$, we choose any of these tuples and find the intersection point $q=\cap_{i=0}^{n}\partial B(p_{i};d_{i})$.
Any other point $p\in(S\setminus\{0,q\})\cap V(\Lambda;0)$ is uniquely determined similarly to the point $q$ above by using its basis distances $d_{0}(p)<d_{1}(p)<\dots<d_{n}(p)$ to points $0=p_{0},p_{1},\dots,p_{n}\in N(\Lambda)$. At the end of reconstruction, we have a final choice between $\pm p$ symmetric with respect to the origin 0. Since the second point $q$ is already fixed, the third point $p$ is also restricted by the distance |pq||p-q| appearing once only in the second and third rows of $\mathrm{PDD}(S;k)$. The distance |pq||p-q| doesn’t help to resolve the ambiguity between $\pm p$ only if $q$ belongs to the bisector of points equidistant to $\pm p$. In this case, $p,0,q$ form a right-angle triangle, which is forbidden by condition (5.6a). Hence $p$ is uniquely determined by the already fixed point $q$ and lattice $\Lambda$.
6 Detecting near-duplicates in the world’s largest databases
This section reports thousands of previously unknown (near-)duplicates in the world’s largest databases [taylor2019million, gravzulis2009crystallography, zagorac2019recent, jain2013commentary]. The sizes in Table 2 below are the numbers of all periodic crystals (with no disorder and full geometric data) in September 2024 (total number is 1,847,462, see Table SM6 and all experimental details in appendix A.
Table 2: Links and versions of the world’s largest materials databases, see their sizes in Table SM6.
database and web address version
CSD: Cambridge Structural Database, http://ccdc.cam.ac.uk version 6.00
COD: Crystallography Open Database, crystallography.net/cod July 30, 2024
ICSD: Inorganic Crystal Structures, icsd.products.fiz-karlsruhe.de Feb 25, 2025
MP: Materials Project, http://next-gen.materialsproject.org v2023.11.1
GNoME: github.com/google-deepmind/materials_discovery Nov 29, 2023
We first used the vector $\mathrm{ADA}(S;100)$ to find nearest neighbors across all databases by $k$-d trees [gieseke2014buffer] up to $L_{\infty}\leq 0.01\textup{\AA }$. Since the smallest inter-atomic distances are about $1\textup{\AA }=10^{-10}$m, atomic displacements up to $0.01\textup{\AA }$ are considered experimental noise. For the closest pairs found by $\mathrm{ADA}(S;100)$, the stronger $\mathrm{PDA}(S;100)$ can have only equal or larger $\mathrm{EMD}\geq L_{\infty}$ by Theorem 4.4. The CSD, COD, ICSD should contain experimental structures. MP is obtained from ICSD by extra optimization.
Table 3 shows that the well-curated 60-year-old CSD has 0.9% near-duplicate crystals, while more than a third of the ICSD consists of near-duplicates that are geometrically almost identical so that all atoms can be matched by an average perturbation up to $0.01\textup{\AA }$. Table 1 in [anosova2024importance, section 6] reported many thousands of exact duplicates, where chemical elements were replaced while keeping all coordinates fixed. These replacements are physically impossible without more substantial perturbations. Five journals are investigating integrity [chawla2023crystallography], see details in appendix A.
The bold numbers in Table 3 count near-duplicates and their percentages within each database, which should be filtered out, else the ground truth data becomes skewed. Table 4 confirms that cell-based comparisons miss near-duplicates as in Fig. 2.
Table 3: Count and percentage of all ideal periodic crystals in each database (left) found to have a near-duplicate in other databases (top) by the distance $\mathrm{EMD}\leq 0.01\textup{\AA }$ on matrices $\mathrm{PDA}(S;100)$.
duplicates CSD COD ICSD MP GNoME
in databases count % count % count % count % count %
CSD 8343 0.92 283000 31.19 26506 2.92 33 0.00 1 0.00
COD 286663 80.18 19568 5.47 47065 13.16 5231 1.46 2705 0.76
ICSD 26853 15.78 69948 41.10 51085 30.01 27194 15.98 15449 9.08
MP 73 0.05 11986 7.82 15188 9.91 19177 12.51 10681 6.97
GNoME 2 0.00 1800 0.47 2614 0.68 3401 0.88 82859 21.53
Table 4: Near-duplicates from Table 3 whose unit cells differ by at least the same threshold of $0.01\textup{\AA }$. Unit cells are compared by $L_{\infty}$ between vectors of corresponding lengths of 3 edges and 3 face diagonals.
duplicates CSD COD ICSD MP GNoME
in databases count % count % count % count % count %
CSD 776 0.09 419 0.05 210 0.02 29 0.00 1 0.00
COD 472 0.13 7263 2.03 8629 2.41 5059 1.42 2684 0.75
ICSD 462 0.27 28863 16.96 42946 25.23 26554 15.60 15360 9.02
MP 70 0.05 11790 7.69 14915 9.73 18582 12.13 10608 6.92
GNoME 2 0.00 1786 0.46 2590 0.67 3346 0.87 60248 15.65
In the past, the (near-)duplicates were impossible to detect at scale, because the traditional comparison through iterative alignment of 15 (by default) molecules by the COMPACK algorithm [chisholm2005compack] is too slow for all-vs-all comparisons. Tables 5 and SM6 compare the running times: minutes of $\mathrm{PDA}(S;100)$ vs years of RMSD, extrapolated for the same machine from the median time 117 milliseconds (582 ms on average) for 500 random pairs in the CSD. On the same 500 pairs, $\mathrm{PDA}(S;100)$ for two crystals and $\mathrm{EMD}$ together took only 7.48 ms on average. All experiments were done on a typical desktop computer (AMD Ryzen 5 5600X 6-core, 32GB RAM).
Table 5: Running times to compute $\mathrm{PDA}(S;100)$ and find all near-duplicates in Table 3 with $\mathrm{EMD}\leq 0.01\textup{\AA }$ across all major databases (seconds in the last 4 columns), compare with years in Table SM6.
database $\mathrm{PDA}$, min:sec $\mathrm{EMD}$, min:sec CSD COD ICSD MP GNoME
CSD 60:44 12:21 125.5 498.1 77.0 19.1 20.6
COD 30:16 16:29 524.5 122.0 235.1 79.6 27.0
ICSD 5:57 22:04 80.5 239.3 515.8 414.9 73.5
MP 1:40 13:31 28.2 82.9 413.8 222.8 63.0
GNoME 4:07 18:59 29.0 26.7 74.5 64.5 943.7
7 Discussion
For hundreds of years, crystals were classified almost exclusively by discrete tools such as space groups or by using reduced cells, which are unique in theory. Fig. 2 (left) showed that any known crystal can be disguised by changing a unit cell, shifting atoms a bit, changing chemical elements, then claimed as ‘new’, see appendix A. Such artificially generated structures threaten the integrity of experimental databases [chawla2023crystallography], which are skewed by previously undetectable near-duplicates. These challenges motivated the stronger questions “how much different?” and “can we get a structure from its code?”, which were formalized in Problem 1.6 aiming for a continuous parametrization of the space of crystals. One limitation is that $\mathrm{PDD}$ is not proved to be complete and a random PDD may not be realizable by a crystal because inter-atomic distances cannot be arbitrary, which we plan to improve in future work for a full solution of Problem 1.6 in the periodic case. However, these invariants already parametrize the ‘universe’ containing all known crystals as ‘shiny stars’ and all not yet discovered crystals hidden in empty spots on the same map. Appendix A shows these geographic-style maps of all four databases in our invariant coordinates.
The impact is the efficient barrier for noisy duplicates of known structures because the invariants quickly find nearest neighbors of newly claimed materials in the existing databases, as shown for all crystals from the GNoME [anosova2024importance] and A-lab [widdowson2025geographic] datasets.
Appendix A Details of experiments on the world’s largest databases
This appendix describes the main experiments in more detail. Some entries in the CSD and COD are incomplete or disordered (not periodic). After removing such entries, we were left with 852,890 CSD structures and 351,380 COD structures.
First we computed $\mu^{(10)}[\mathrm{PDD}(S;100)]$ for all entries, taking 27 min 33 sec for the CSD and 12 mins 15 sec for COD (2 ms per structure on average). To find exact geometric matches between databases, we use the $k$-d tree data structure, designed for fast nearest neighbor lookup. A $k$-d tree can be constructed from any collection of vectors, which can then be queried for a number of nearest neighbors of a new vector, using a binary tree style algorithm with logarithmic search time.
Then we flattened each matrix $\mu^{(10)}[\mathrm{PDD}(S;100)]$ to a vector with 1000 dimensions, constructed a $k$-d tree for both CSD and COD, then queried the 10 nearest neighbors for each item in the other. If the most distant neighbor for any entry is closer than the threshold $10^{-10}\textup{\AA }$ (within floating point error), we extend the search and find more neighbors until all pairs within the threshold are found. We found a total of 278,236 geometric matches (almost exact duplicates at the atomic level); an overlap between the databases of one third of the CSD and over 80% of the COD.
Of particular interest are the 235 pairs with near-zero distance but different chemical compositions. Indeed, the impossibility of complex organic structures sharing the exact same geometry but not composition implies an error or labeling issue. All the pairs were confirmed as geometric duplicates by manually checking their CIFs and found to have different compositions, mostly for the three reasons given below. The 5 remaining pairs not in these three categories are in Table 9 below.
  • The source CIF has atoms whose types are labelled differently by the tags ‘_atom_site_label’ and ‘_atom_site_type_symbol’. COD entries always use the data in the uploaded CIF, but CSD entries occasionally have data corrected and if so often have a remark describing the correction (109 pairs, Table 6).
  • Disorder was modeled as a ‘mixed site’ with one atomic type present and a remark on the CSD entry explaining the disorder (20 pairs, Table 8).
  • Types in the CIF are consistent but CSD curators discovered incorrectly labelled atoms which were corrected and given a remark (78 pairs, Table 7).
CSD ID COD ID
ABAGUG 4112689
AFUXEG 2238369
AJAREI 7113511
AJAREI 7103824
BAKXUH 8100721
BIGNUA 5000340
BOQBAK 2009202
CABSAA 2200584
CALWIW 4114997
CAQFUV 7027367
CUDJAP 1557108
DECJUS 4065161
DECTAI 4065524
DEGFOL 2208310
DEHKUX 7101047
DOBBIF 7213201
DUDZOS 4302088
EBASIN01 7708085
EFESUE 4107864
EGELUY 4108535
ELOJOE 4314231
ENIZEH 2018012
ESADAD 4062269
EVEMIB 4020894
EXATEC 7050257
EXATIG 7050258
FONGAQ01 2005101
FUPJIJ 7212965
GESJIY 4333010
GETSAD 7245388
GUHYOX 7010289
HABTAF 2001740
HIXWEQ 2008462
IKOSIL 4065905
JECBID 7006569
JUCJOJ 4003435
KABHOL 4113866
CSD ID COD ID
KAVYOW 7008840
KEBQUF 7018464
KEZLOS 4117778
KEZMUZ 4117772
KIZFOR 7232188
KIZJOT 4029575
LABSAI 2001822
LAMQEV 4116446
LAVFAP 2001334
LAZWOY 2009422
LINLOJ 2003397
LINLUP 2003398
LUNDIH 1507498
MEHCEI 2208583
MEJRAV 4101504
MENCAJ 7009977
METSAF 7702634
NAJQUK 4323901
NEDXID01 2105611
NIQJIJ01 1549188
NOCXIM 4322709
NOVHUB02 2103787
NUMWOH 2007448
NUVZOV 4501471
ODEBII 4115837
OGOLUR 5000295
OHEFAI 7012100
OHEJIU 7204467
PAMWIK 2205526
PAXKEG 2235126
PAYSUF 2235091
PHOXBZ01 2017696
PIHJUL 4030494
QAHFOV 7012335
QAZTEQ 4077596
QEJYUA 4508631
CSD ID COD ID
QIQNIN 4077174
QOQFOT 4348248
QUXBAN 2017697
RAKMOF 7114739
RARFUM 4327332
RIVKOW01 4310386
ROCJUP 4304894
RORGUA 4323669
RORGUA02 4323669
RUVFET 4323710
SAQHIC 1100776
SAQQUX 4308912
SAXCUP 2007898
TIPYOG 2005914
TOCNOO 4323981
UJECOB 7012760
UJIKAZ 7213431
UVOHIY 7040448
WASKAC 2001382
WATMIO 4309447
WIKRIS 8102105
WIRJEM 2005120
XAFDUD 4321242
XAGJUK 8101251
XAVDEF 4103386
XIHVOZ 4317724
XIJNOT 4115818
XOFXIZ 1507458
XOFXOF 1507459
XOPNAT 7218637
XUFLUH 7034643
YEJQAF 2012123
ZAGCUJ 1559337
ZAYRUM 2003941
ZEXQUO 2004127
ZIKMAH 2004275
Table 6: 109 exact geometric matches (within $10^{-10}\textup{\AA }$) between the CSD and COD where the original CIF has atoms labelled as different types by ‘_atom_site_label’ and ‘_atom_site_type_symbol’. Several of the CSD entries have a remark noting that atoms were corrected in curation.
CSD ID COD ID
AFUKIX 7211182
AJIRAM01 2100097
BAPLOT09 7121265
BASLAJ 7050473
BASMAK 7050478
BEPWUQ 4507409
BIHVUL 7210243
BIKJEN 7231097
BODZEB 7215818
BOMMEX 4124237
CIPDIQ 7213596
COLNUP 4034420
COTNAC 7219615
DAGRUB 4349194
DENBAD 7710591
DIBGAX 7151087
DISNAW 1543965
DOJFEQ 7230639
DOSSOW 7123961
DUFXOS 7104457
EMUMEF 1503106
ETEPIC 2203286
EWABIO 4324780
FEBBOH 7130024
FOBXAY 7122779
FORWOA 7116555
GACZEQ 7151378
IQAFEN 7225754
ISORIU 7242793
ISUFAE 7205743
JEMLAP 4101489
JOHXUB 7114582
KUKQIS 7234247
CSD ID COD ID
KUTWUU 7126770
LEBTET 7110143
MARSIH 4321045
MAVZUG 7107511
MIPNEG 4335723
MOGHAU 7123768
NAFTIA 7223916
NAJMER 7050031
NEHFUE 4131268
NEYJIM 7021415
NIFJAO 4022923
NIMXOY 4334458
NUKCAZ 7035092
NUQVAY 7118051
OKUJOV 4347519
OMIJIF 7118994
PAQCEQ 4061419
PECRUL 4300654
PIBTAW 1505325
PICFIR 4072624
PIGJEW 4080504
PINHUP 1558382
PUTCOY 7055058
QAMKAU 7705818
QANLIE 7061176
QOTVUS 7221578
QOWKOE 4341138
QUCXAP 7117360
RADBAB 7025360
REGVII 4116980
REMVOU 2006347
REYRES 4116989
CSD ID COD ID
RIDYAI 7131471
ROBKID01 1520266
SALGUK 7155485
SELHAU 4027023
SIJBAQ 7109679
SOVZOT 4063498
TAVWEW 7129345
TEMMOQ 7056766
TETQUI 7711227
UCACAF 7119310
UGOVER 4115188
UGUBIJ 7220063
UGUSIB 1551384
UKAXUB 7234657
UMESIQ 7225104
UVOHOE 7040449
UYEBUX 7236357
VEFLUR 1561274
VENJIJ 4331164
VOCNUY 7239443
WOTMEA 4036052
WULGIV 4036188
XEXCOV 7045895
YEJNOU 7710456
YEPSUI 8000091
YURCEV 7036965
YUYDAZ 7037146
ZEYKIA 7230274
ZIDBOF 7210579
ZIGDIG 7246585
ZUNNUU 7059654
ZITXUV 2004330
Table 7: 97 exact geometric matches (within floating point error of $10^{-10}\textup{\AA }$) between the CSD and COD with different chemical compositions where erroneously labelled atoms were corrected by the CSD entry in curation. Most entries have a remark mentioning the correction.
CSD ID COD ID
FIQDUI 7713232
GODSEY 4305065
GOHPAU01 2102515
LIJXAD 2010401
LIJXEH 2010402
MUMXIB01 2102385
NUTZIN 7036504
CSD ID COD ID
NUTZOT 7036505
QALLUL 4505437
TIPSAM 2101647
TIPSAM01 2101646
TIPSAM02 2101648
TOGVOA 2005985
ZUGVIG 2004797
CSD ID COD ID
ZUGVOM 2004798
ZUGVUS 2004799
ZUGWAZ 2004800
ZUGWED 2004801
ZUGWIH 2004802
ZUHCOW 2004740
Table 8: 20 exact geometric matches (within floating point error $10^{-10}\textup{\AA }$) between the CSD and COD with different compositions where disorder was modeled as a ‘mixed site’ with only one of two atomic types listed. Usually the CSD entry has a remark describing the disorder.
CSD ID COD ID Remark
APEJUD 1544509 APEJUD has atom label ‘Unknown1?’
HIWHEA 4321802 C1 $\leftrightarrow$ N1C
IPOQOU 4063641 N2 $\leftrightarrow$ C22
LEFYIF 4300748 B1, B2, C5, C1 $\leftrightarrow$ C27, C17, B21, B11
NIDPIB 7208250 N2 $\leftrightarrow$ O21
Table 9: 5 exact geometric matches (within $10^{-10}\textup{\AA }$) between the CSD and COD with different compositions. It could not be confirmed if the last four pairs are erroneous or corrected by the CSD.
In addition to cross-comparing the CSD and COD, we also analyzed the ICSD and Materials Project database (MP) and compared them all pairwise, as well as searching for duplicates within each database. Table 10 below shows how many matches were found, and how many also shared the same composition.
databases matches same composition
CSD vs COD 276,494 276,376
CSD vs ICSD 3,272 3,270
COD vs ICSD 35,162 32,023
COD vs MP 14 4
ICSD vs MP 71 32
Table 10: Number of exact matches ($\mathrm{EMD}$ within $10^{-10}\textup{\AA }$) between the four major databases.
Table 11: These times for all comparisons by COMPACK [chisholm2005compack] are extrapolated on the same machine, which completed Table 3 of near-duplicates across all the major databases within 2 hours.
database periodic crystals unordered pairs COMPACK time, sec years
CSD 907,246 411,547,198,635 $4.81\times 10^{10}$ 1526
COD 357,510 63,906,521,295 $7.48\times 10^{9}$ 237
ICSD 170,206 14,484,956,115 $1.69\times 10^{9}$ 53
MP 153,235 11,740,405,995 $1.37\times 10^{9}$ 43
GNoME 384,938 74,088,439,453 $8.67\times 10^{9}$ 274
Figure 4: The projections of the CSD in the invariants $\mathrm{PPC},\mathrm{ADA}_{1},\mathrm{ADA}_{2},\mathrm{ADA}_{3}$.
Figure 5: The projections of the COD in the invariants $\mathrm{PPC},\mathrm{ADA}_{1},\mathrm{ADA}_{2},\mathrm{ADA}_{3}$.
Figure 6: The projections of the ICSD in the invariants $\mathrm{PPC},\mathrm{ADA}_{1},\mathrm{ADA}_{2},\mathrm{ADA}_{3}$.
Figure 7: The projections of the MP in the invariants $\mathrm{PPC},\mathrm{ADA}_{1},\mathrm{ADA}_{2},\mathrm{ADA}_{3}$.
Table 12 compares the proven properties of past and new descriptors.
Descriptor Invariant Continuity Complete Reconstruction Time
primitive cell $\times$ $\times$ $\times$ $\times$
reduced cell $\times$ $\times$ $\times$
space group $\times$ $\times$ $\times$
PDF [terban2021structural] $\times$ $\times$ ✓*
MACE [batatia2022mace] $\times$ ✓* $\times$ ✓*
densities [edelsbrunner2021density] ✓* $\times$ ✓*
AMD [widdowson2022average] $\times$ $\times$
PDD [widdowson2022resolving] ✓* ✓*
isosets [anosova2021isometry, anosova2026recognition] ✓*
Table 12: Comparison of crystal descriptors in the context of Problem 1.6. ✓* in the ‘Computable’ column indicates that only an approximate algorithm exists for distances, and ✓* in the ‘Complete’ and ‘Reconstruction’ columns means that the condition holds in general position. For example, all invariants based on local atomic environments, such as MACE [batatia2022mace], discontinuously change under almost any perturbation that arbitrarily scales up a primitive cell as in Fig. 2 (left), except the complete isosets [anosova2021isometry] whose continuous metric was designed to be Lipschitz continuous [anosova2026recognition].
Appendix B Examples and instructions for the PDD code and data
This appendix explains the code at https://pypi.org/project/average-minimum-distance.
B.1 Pseudocode for computing Pointwise Distance Distributions
The algorithm accepts any periodic point set $S\subset\mathbb{R}^{n}$ in the form of a unit cell $U$ and a motif $M\subset S$. The cell is given as a square $n\times n$ matrix with basis vectors in the columns, and the motif points in Cartesian form lying inside the unit cell. For dimension 3, the typical Crystallographic Information File (CIF) with six unit cell parameters and motif points in terms of the cell basis is easily converted to this format. Otherwise, the unit cell and motif points can be given directly, in any dimension. Specifically, the PDD function’s interface is as follows:
Input:
  • motif: array shape $(m,n)$. Coordinates of motif points in Cartesian form.
  • cell: array shape $(n,n)$. Represents the unit cell in Cartesian form.
  • k: int >0>0. Number of columns to return in PDD(S;k)(S;k).
Output:
  • pdd: array with $k+1$ columns.
Before giving the pseudocode, we outline the key objects and functions in use:
  • A generator g, which creates points from the set $S$ to find distances to,
  • KDTrees (canonically $k$ is the dimension here, in our case it’s denoted $n$), data structures designed for fast nearest-neighbor lookup in $\mathbb{R}^{n}$.
Once g is constructed, next(g) is called to get new points from the infinite set $S$. The first call returns all points in the given unit cell (i.e. the motif), and successive calls returns points from unit cells further from the origin in a spherical fashion.
A KDTree is constructed with a point set $T$, then queried with another $Q$, returning a matrix with distances from all points in $Q$ to their nearest neighbors (up to some given number, $k$ below) in $T$, as well as the indices of these neighbors in $T$.
The functions collapse_equal_rows and lexsort_rows, which perform the collapsing and lexicographical sorting steps of computing PDD, respectively, are assumed to be implemented elsewhere. The following pseudocode finds $\mathrm{PDD}(S;k)$ for a periodic set $S$ described by motif and cell:
def PDD(motif, cell, k):
    cloud = [] # contains points from S
    g = point_generator(motif, cell)
    # at least k points will be needed
    while len(cloud) < k:
        points = next(g)
        cloud.extend(points)
    # first distance query
    tree = KDTree(cloud)
    D_, inds = tree.query(motif, k)
    D = zeros_like(D_)
    # repeat until distances don’t change,
    # then all nearest neighbors are found
    while not D == D_:
        D = D_
        cloud.extend(next(g))
        tree = KDTree(cloud)
        D_, inds = tree.query(motif, k)
    pdd = collapse_equal_rows(D_)
    pdd = lexsort_rows(pdd)
    return pdd
B.2 Instructions for the attached PDD code and specific examples
A Python script implementing Pointwise Distance Distributions along with examples can be found in the zip archive included in this submission. Python 3.7 or greater is required. The dependency packages are NumPy (<1.22<1.22), SciPy ($\geq 1.6.1$), numba ($\geq 0.55.0$) and ase ($\geq 3.22.0$); if you do not wish to affect any currently installed versions on your machine, create and activate a virtual environment before the following.
Unzip the archive and in a terminal navigate to the unzipped folder. Install the requirements by running pip install -r requirements.txt. Run python followed by the example script of choice, and then any arguments (outlined below), e.g.
    \$ python kite_trapezium_example.py
    trapezium: [(0, 0), (1, 1), (3, 1), (4, 0)]
    PDD:
    [[0.5        1.41421356 2.         3.16227766]
     [0.5        1.41421356 3.16227766 4.        ]]
    kite: [(0, 0), (1, 1), (1, -1), (4, 0)]
    PDD:
    [[0.25       1.41421356 1.41421356 4.        ]
     [0.5        1.41421356 2.         3.16227766]
     [0.25       3.16227766 3.16227766 4.        ]]
    EMD between trapezium and kite: 0.874032
Here is the list of included example scripts and their parameters:
  • kite_trapezium_example.py prints the PDDs of the 4-point sets $K$ (kite) and $T$ (trapezium) in Fig. 8 (left), along with their EMD.
    Figure 8: Left: the 4-point sets $K=\{(\pm 2,0),\;(\pm 1,1)\}$ and $T=\{(\pm 2,0),(-1,\pm 1)\}$ have the same pairwise distances $\sqrt{2},\sqrt{2},2,\sqrt{10},\sqrt{10},4$. Right: the sequences $S(r)=\{0,r,2+r,4\}+8\mathbb{Z}$ and $Q(r)=\{0,2+r,4,4+r\}+8\mathbb{Z}$ for $0<r\leq 1$ have the same Patterson function [patterson1944ambiguities, p. 197, Fig. 2].
  • 1D_sets_example.py shows that the 1D periodic sets in Fig. 8 (right) are distinguished by their PDDs for any $0<r\leq 1$. This script requires $r$ to be passed after the file name, e.g. ‘python 1D_sets_example.py 0.5’.
  • T2_14_15_example.py compares the crystals shown in Fig. 9, whose original CIFs are included. This optionally accepts the number $k$ of columns in the computed PDD, e.g. ‘python T2_14_15_example.py --k 50’ compares by PDD with $k=50$. If not included, $k=100$ is used as the default.
    Figure 9: Crystals 14, 15 from the database of 5679 simulated crystals reported in [pulido2017functional] consist of identical T2 molecules and have very different Crystallographic Information Files (with different motifs in unit cells of distinct shapes) but are nearly identical under isometry.
  • CSD_duplicates_example.py computes and compares the PDDs of isometric crystals from the CSD discussed in section A, giving distances of exactly zero. This optionally accepts the parameter $k$ controlling the number of columns in the computed PDD, in the same way as T2_14_15_example.py.
If you wish to run the code on your own sets or CIF files, you can use the functions exposed in the main script pdd.py. Use pdd.read_cif() to parse a cif and return a crystal, or define one manually as a tuple (motif, cell) with NumPy arrays. Pass this as the first argument to pdd.pdd() with an integer k as the second to compute the PDD. Pass two PDDs to pdd.emd() to calculate the Earth mover’s distance between them. For finite sets, the function pdd.pdd_finite() accepts just one argument, an array containing the points, and returns the PDD. Figures 10, 11, 12, 13, 14, show near-duplicates with very different cells, which were counted in Table 4.
Figure 10: In the CSD, near-duplicates PUBTEM (left) and PUBTEM01 (right) have a very small $\mathrm{EMD}=0.00038\textup{\AA }$ on invariants $\mathrm{PDA}(S;100)$, though their unit cells are rather different.
Figure 11: In the COD, near-duplicates 2310812 (top) and 2310813 (bottom) have a very small $\mathrm{EMD}=0.0008\textup{\AA }$ on invariants $\mathrm{PDA}(S;100)$, though their unit cells differ by a factor of about 3.
Figure 12: In the ICSD, near-duplicates 42291 (top) and 42302 (bottom) have a very small $\mathrm{EMD}=0.0024\textup{\AA }$ on invariants $\mathrm{PDA}(S;100)$, though their unit cells differ by a factor of about 3.
Figure 13: In the MP, near-duplicate entries mp-90 (left) and mp-1221808 (right) have a very small $\mathrm{EMD}=0.0087\textup{\AA }$ on invariants $\mathrm{PDA}(S;100)$, though their unit cells substantially differ.
Figure 14: In the GNoME, near-duplicates 4cb3b6ed9f (left) and 776c1b7570 (right) in the GNoME have $\mathrm{EMD}=0.0079\textup{\AA }$ on invariants $\mathrm{PDA}(S;100)$, though their unit cells are very different.
Appendix C Detailed proofs of auxiliary lemmas and Theorem 4.2
This appendix proves Lemmas 3.4-3.5, which were used in Theorem 3.6, and Theorem 4.2.
Proof of Lemma 3.4.
Intersect the three regions ${U^{-}}(p;r)\subset C(p;r)\subset{U^{+}}(p;r)$ with $S$ in $\mathbb{R}^{n}$ and count all points: $|S\cap{U^{-}}(p;r)|\leq|S\cap C(p;r)|\leq|S\cap{U^{+}}(p;r)|$.
The union ${U^{-}}(p;r)$ consists of $m^{-}(p;r)=\dfrac{\mathrm{vol}[U^{-}(p;r)\cap R^{l}]}{\mathrm{vol}[U]}$ shifted cells, which all have the same volume $\mathrm{vol}[U\cap R^{l}]$. Since $|S\cap U|=m$, we get $|S\cap U^{-}(p;r)|=\dfrac{\mathrm{vol}[U^{-}(p;r)\cap R^{l}]}{\mathrm{vol}[U]}m$. Similarly, we count all points of $S$ in the upper union as follows: $|S\cap{U^{+}}(p;r)|=\dfrac{\mathrm{vol}[{U^{+}}(p;r)\cap R^{l}]}{\mathrm{vol}[U]}m$. The bounds for $|S\cap C(p;r)|$ become
$\dfrac{\mathrm{vol}[{U^{-}}(p;r)\cap R^{l}]}{\mathrm{vol}[U]}m\leq|S\cap C(p;r)|\leq\dfrac{\mathrm{vol}[{U^{+}}(p;r)\cap R^{l}]}{\mathrm{vol}[U]}m,$
which proves the internal inequalities $m^{-}(p;r)m\leq|S\cap C(p;r)|\leq m^{+}(p;r)m$. Then
$\mathrm{vol}[{U^{-}}(p;r)\cap R^{l}]\leq\dfrac{\mathrm{vol}[U\cap R^{l}]}{m}|S\cap C(p;r)|\leq\mathrm{vol}[{U^{+}}(p;r)\cap R^{l}].$
For the width $w$ of the unit cell $U$, the smaller cylinder C(p;rw)C(p;r-w) is completely contained within the lower union ${U^{-}}(p;r)$. Indeed, if $|\vec{q}-\vec{p}|\leq r-w$, then $q\in U+\vec{v}$ for some $\vec{v}\in\Lambda$. Then $(U+\vec{v})$ is covered by the cylinder $C(q;w)$, hence by $C(p;r)$ due to the triangle inequality. The inclusion $C(p;r-w)\subset{U^{-}}(p;r)$ implies the lower bound for the volumes: $(r-w)^{l}V_{l}=\mathrm{vol}[C(p;r-w)\cap R^{l}]\leq\mathrm{vol}[{U^{-}}(p;r)\cap R^{l}]$, where $V_{l}$ is the unit ball volume in $\mathbb{R}^{l}$. Then $\dfrac{(r-w)^{l}V_{l}}{\mathrm{vol}[U\cap R^{l}]}\leq\dfrac{\mathrm{vol}[{U^{-}}(p;r)\cap R^{l}]}{\mathrm{vol}[U\cap R^{l}]}=m^{-}(p;r)$, which implies the first required inequality in the lemma:
$\left(\dfrac{r-w}{\mathrm{PPC}(S)}\right)^{l}=\dfrac{(r-w)^{l}mV_{l}}{\mathrm{vol}[U\cap R^{l}]}\leq\dfrac{\mathrm{vol}[{U^{-}}(p;r)\cap R^{l}]}{\mathrm{vol}[U\cap R^{l}]}m=m^{-}(p;r)m.$
The last required inequality is proved similarly by using $U^{+}(p;r)\subset C(p;r+w)$.
Proof of Lemma 3.5.
Let $q\in S$ be a $k$-th neighbor of $p$ in $S$. There can be several points $q\in S$ at the distance $|q-p|=d_{k}(S;p)$ but the argument below works for any $q$. The closed cylinder $C(p;r)$ with $r=d_{k}(S;p)$ contains the $k$-th neighbor $q$ of $p$ and hence has more than $k$ points (including $p$) from $S$. The upper bound of Lemma 3.4 for $r=d_{k}(S;p)$ implies that $k<|S\cap C(p;r)|\leq\dfrac{(r+w)^{l}}{(\mathrm{PPC}(S))^{l}}$. Taking the $l$-th roots gives $\sqrt[l]{k}<\dfrac{r+w}{\mathrm{PPC}(S)}$, so $r=d_{k}(S;p)>\mathrm{PPC}(S)\sqrt[l]{k}-w$.
For any radius $r$ such that $\sqrt{r^{2}+h^{2}}<d_{k}(S;p)$, the closed cylinder $C(p;r)$ contains only points at a maximum distance $\sqrt{r^{2}+h^{2}}$ from $p$. Then $C(p;r)$ does not include the $k$-th neighbor $q$ of $p$ and hence contains at most $k$ points (including $p$) from $S$. The lower bound of Lemma 3.4 for $r<\sqrt{(d_{k}(S;p))^{2}-h^{2}}$ implies that $\dfrac{(r-w)^{l}}{(\mathrm{PPC}(S))^{l}}\leq|S\cap C(p;r)|\leq k$. Since the inequality $\dfrac{(r-w)^{l}}{(\mathrm{PPC}(S))^{l}}\leq k$ holds for the constant upper bound $k$ and any radius $r<\sqrt{(d_{k}(S;p))^{2}-h^{2}}$, the same inequality holds for the radius $r=\sqrt{(d_{k}(S;p))^{2}-h^{2}}$. Then $\dfrac{r-w}{\mathrm{PPC}(S)}\leq\sqrt[l]{k}$,
$r=\sqrt{(d_{k}(S;p))^{2}-h^{2}}\leq\mathrm{PPC}(S)\sqrt[l]{k}+w,\quad d_{k}(S;p)\leq\sqrt{(\mathrm{PPC}(S)\sqrt[l]{k}+w)^{2}+h^{2}}.$
Example C.1 (stronger asymptotic $\mathrm{ADA}_{k}(S)\to 0$ as $k\to+\infty$ for $\mathbb{Z}^{n}$).
The survey [ivic2004lattice] describes progress on the generalized Gauss circle problem expressing the number of points from the cubic lattice $\mathbb{Z}^{n}$ within a ball of a radius $r$ as $k=V_{n}r^{n}-O(r^{\alpha_{n}+\varepsilon})$ for any $\varepsilon>0$, where $\alpha_{n}<n-1$ for $n\geq 2$. The cubic lattice has $\mathrm{PPC}(\mathbb{Z}^{n})=1/\sqrt[n]{V_{n}}$. Let $d_{k}$ denote the distance from the origin 0 to its $k$-th neighbor in $\mathbb{Z}^{n}$. Then $k=V_{n}d_{k}^{n}-O(d_{k}^{\alpha_{n}+\varepsilon})$, so $d_{k}=\sqrt[n]{\dfrac{k+O(d_{k}^{\alpha_{n}+\varepsilon})}{V_{n}}}=\mathrm{PPC}(\mathbb{Z}^{n})\sqrt[n]{k+O(d_{k}^{\alpha_{n}+\varepsilon})}$. Then
$\dfrac{\mathrm{ADA}_{k}(\mathbb{Z}^{n})}{\mathrm{PPC}(\mathbb{Z}^{n})}=\dfrac{d_{k}}{\mathrm{PPC}(\mathbb{Z}^{n})}-\sqrt[n]{k}=\sqrt[n]{k+O(d_{k}^{\alpha_{n}+\varepsilon})}-\sqrt[n]{k}=\dfrac{O(d_{k}^{\alpha_{n}+\varepsilon})}{P_{n}(\sqrt[n]{k+O(d_{k}^{\alpha_{n}+\varepsilon})},\sqrt[n]{k})},$
where $P_{n}$ is a homogeneous polynomial of degree n1n-1, e.g. $P_{2}(x,y)=x+y$, $P_{3}(x,y)=x^{2}+xy+y^{2}$. Since the numerator has the power $\alpha_{n}<n-1$ of $d_{k}=O(\sqrt[n]{k})$ for $n\geq 2$, the final expression and hence $\mathrm{ADA}_{k}(\mathbb{Z}^{n})$ have limit 0 as $k\to+\infty$.
Theorem 4.1 will be proved similar to [widdowson2022average, Theorem 13] by Lemmas C.2, C.3, C.4. Partial cases of Lemmas C.2 and C.3 appeared for $l=n$ in [edelsbrunner2021density, Lemma 2] and for $\mathbb{R}^{n}$ in [widdowson2022average, Lemma 8], respectively.
Lemma C.2 (common lattice).
Let $l$-periodic point sets $S,Q\subset\mathbb{R}^{n}$ have a bottleneck distance $d_{B}(S,Q)<\min\{r(S),r(Q)\}$. Then $S,Q$ have a common lattice $\Lambda$ with a unit cell $U$ such that $S=\Lambda+(U\cap S)$ and $Q=\Lambda+(U\cap Q)$.
Proof of Lemma C.2.
Choose the origin $0\in\mathbb{R}^{n}$ at a point of $S$. Applying translations, we can assume that primitive unit cells $U(S),U(Q)$ of the given $l$-periodic sets $S,Q$ have a vertex at the origin 0. Then $S=\Lambda(S)+(U(S)\cap S)$ and $Q=\Lambda(Q)+(U(Q)\cap Q)$, where $\Lambda(S),\Lambda(Q)$ are $l$-dimensional lattices of $S,Q$, respectively. We are given that every point of $Q$ is $d_{B}(S,Q)$-close to a point of $S$, where the bottleneck distance $d_{B}(S,Q)$ is strictly less than the packing radius $r(Q)$.
Assume by contradiction that $S,Q$ have no common lattice. Then there is a point $p\in\Lambda(S)\subset S$ whose all integer multiples $k\vec{p}\in\Lambda(S)$ do not belong to $\Lambda(Q)$ for $k\in\mathbb{Z}-\{0\}$. Any such multiple $k\vec{p}\in\Lambda(S)\subset S$ can be translated by a vector of $\Lambda(Q)$ to a point $t(k)$ in the unit cell $U(Q)$ so that $k\vec{p}\equiv t(k)\pmod{\Lambda(Q)}$. Since the cell $U(Q)$ contains infinitely many points $t(k)$ for $k\neq 0$, one can find a pair $t(i)\neq t(j)$ at a distance less than $\delta=r(Q)-d_{B}(S,Q)>0$. For any $m\in\mathbb{Z}$, the following points are equivalent modulo (translations along the vectors of) the lattice $\Lambda(Q)$.
$t(i+m(j-i))\equiv(i+m(j-i))\vec{p}=i\vec{p}+m(j\vec{p}-i\vec{p})\equiv t(i)+m(t(j)-t(i)).$
These points for $m\in\mathbb{Z}$ lie in a straight line with gaps $|t(j)-t(i)|<\delta$. The open balls with the packing radius $r(Q)$ and centers at all points of $Q$ do not overlap. Hence all closed balls with the radius $d_{B}(S,Q)<r(Q)$ and the same centers are at least $2\delta$ away from each other. Due to $|t(j)-t(i)|<\delta=r(Q)-d_{B}(S,Q)$, there is $m\in\mathbb{Z}$ such that t(i)+m(t(j)t(i))t(i)+m(t(j)-t(i)) is outside the union $Q+\bar{B}(0;d_{B}(S,Q))$ of all these smaller balls. Then t(i)+m(t(j)t(i))t(i)+m(t(j)-t(i)) has a distance more than $d_{B}(S,Q)$ from any point of $Q$. The translations along all vectors of the lattice $\Lambda(Q)$ preserve the union of balls $Q+\bar{B}(0;d_{B}(S,Q))$. Then the point $(i+m(j-i))\vec{p}\in\Lambda(S)\subset S$, which is equivalent to t(i)+m(t(j)t(i))t(i)+m(t(j)-t(i)) modulo $\Lambda(Q)$, has a distance more than $d_{B}(S,Q)$ from any point of $Q$. This conclusion contradicts the definition of $d_{B}(S,Q)$.
Lemma C.3 (perturbed distances).
For some $\varepsilon>0$, let $g:S\to Q$ be a bijection between any discrete sets in a space $X$ with a metric $d_{X}$ such that $d_{X}(g(p),p)\leq\varepsilon$ for all $p\in S$. Then, for any $i\geq 1$, let $p_{i}\in S$, $\tilde{p}_{i}\in Q$ be the $i$-th nearest neighbors of $p\in S$, $\tilde{p}=g(p)\in Q$, respectively. Then the distances from the points $p,\tilde{p}$ to their $i$-th neighbors $p_{i},\tilde{p}_{i}$ in $X$ are $2\varepsilon$-close to each other, i.e. $|d_{X}(p,p_{i})-d_{X}(\tilde{p},\tilde{p}_{i})|\leq 2\varepsilon$.
Proof of Lemma C.3.
Shifting the point $g(p)$ back to $p$, assume that $p=g(p)$ is fixed and all other points change their positions by at most $2\varepsilon$. Assume by contradiction that the distance from $p$ to its new $i$-th neighbor $t_{i}$ is less than $d_{X}(p,p_{i})-2\varepsilon$. Then all first new $i$ neighbors $\tilde{p}_{1},\dots,\tilde{p}_{i}\in Q$ of $p$ belong to the open ball with the center $p$ and radius $d_{X}(p,p_{i})-2\varepsilon$. Since the bijection $g$ shifted every $\tilde{p}_{1},\dots,\tilde{p}_{i}$ by at most $2\varepsilon$, their preimages $g^{-1}(\tilde{p}_{1}),\dots,g^{-1}(\tilde{p}_{i})$ belong to the open ball with the center $p$ and the radius $d_{X}(p,p_{i})$. Then the $i$-th neighbor of $p$ within $S$ is among these $i$ preimages, i.e. the distance from $p$ to its $i$-th nearest neighbor should be strictly less than the assumed value $d_{X}(p,p_{i})$. We similarly get a contradiction by assuming that the distance from $p$ to its new $i$-th neighbor $\tilde{p}_{i}$ is more than $d_{X}(p,p_{i})+2\varepsilon$.
Lemma C.4 (perturbed distance vectors).
For $\varepsilon>0$, let $g:S\to Q$ be a bijection between any discrete sets in a space $X$ with a metric $d_{X}$ so that $d_{X}(g(p),p)\leq\varepsilon$ for all $p\in S$. Then $g$ changes the vector $\vec{R}(S,p)=(d_{X}(p,p_{1}),\dots,d_{X}(p,p_{k}))$ of the first $k$ minimum distances from any point $p\in S$ to its $k$ nearest neighbors $p_{1},\dots,p_{k}\in S$ by at most $2\varepsilon\sqrt[q]{k}$ in the distance $L_{q}$. So if $\vec{R}(Q,\tilde{p})=(d_{X}(\tilde{p},\tilde{p}_{1}),\dots,d_{X}(\tilde{p},\tilde{p}_{k}))$ is the vector of the first $k$ minimum distances from $\tilde{p}=g(p)$ to its $k$ nearest neighbors $\tilde{p}_{1},\dots,\tilde{p}_{k}$ in $Q$, then $L_{q}(\vec{R}(S,p),\vec{R}(Q,\tilde{p}))\leq 2\varepsilon\sqrt[q]{k}$.
Proof of Lemma C.4.
By Lemma C.3, every coordinate of $\vec{R}(S,p)$ changes by at most $2\varepsilon$. Hence the distance $L_{q}(\vec{R}(S,p),\vec{R}(Q,\tilde{p}))\leq\big(\sum\limits_{i=1}^{k}(2\varepsilon)^{q}\big)^{1/q}=2\varepsilon\sqrt[q]{k}$.
Proof of Theorem 4.2.
The bottleneck distance between the given sets $S,Q\subset X$ is $d_{B}(S,Q)=\inf\limits_{g:S\to Q}\;\sup\limits_{p\in S}d_{X}(g(p),p)$. Then for any $\delta>0$ there is a bijection $g:S\to Q$ such that $\sup\limits_{p\in S}d_{X}(g(p),p)\leq d_{B}(S,Q)+\delta$. If the given sets $S,Q$ are finite, one can set $\delta=0$. Indeed, there are only finitely many bijections $g:S\to Q$, hence the infimum in the definition above is achieved for one of these bijection $g$.
(a) For any discrete sets $S,Q\subset X$ be with finite subsets $M,T$ of the same number $m$ of points, respectively, we use the notations of Definition 3.1. The given 1-1 perturbation $g:S\to Q$ defines the simplest 1-1 flow from the row of any $p\in M$ in the matrix $D(S,M;k)$ to the row of $g(p)\in T$ in $D(Q,T;k)$ by setting $f_{ii}=\frac{1}{m}$ and $f_{ij}=0$ for $i\neq j$, where $i,j=1,\dots,m$. All rows of $D(S,M;k)$ that are identical to each other are collapsed to a single row, similarly for $D(Q,T;k)$. By summing up weights of all collapsed rows, the above flow induces a flow from all distance vectors in $\mathrm{PDD}(S,M;k)$ to all distance vectors in $\mathrm{PDD}(Q,T;k)$.
Then $\mathrm{EMD}_{q}(\mathrm{PDD}(S,M;k),\mathrm{PDD}(Q,T;k))\leq\frac{1}{m}\sum\limits_{i=1}^{m}L_{q}(\vec{R}_{i}(S),\vec{R}_{i}(Q))$, because $\mathrm{EMD}_{q}$ minimizes the cost in Definition 4.2. The upper bound $L_{q}(\vec{R}_{i}(S),\vec{R}_{i}(Q))\leq 2(\varepsilon+\delta)\sqrt[q]{k}$ from Lemma C.4 implies that
$\mathrm{EMD}_{q}(\mathrm{PDD}(S,M;k),\mathrm{PDD}(Q,T;k))\leq\frac{1}{m}\sum\limits_{i=1}^{m}2(\varepsilon+\delta)\sqrt[q]{k}=2(\varepsilon+\delta)\sqrt[q]{k},$
which holds for any small $\delta>0$. By taking the limit for $\delta\to 0$, we get the required upper bound $\mathrm{EMD}_{q}(\mathrm{PDD}(S,M;k),\mathrm{PDD}(Q,T;k))\leq 2\varepsilon\sqrt[q]{k}$.
(b) In the $l$-periodic case by Lemma C.2, the given sets $S,Q$ should have a common $l$-dimensional lattice $\Lambda$. Any primitive cell $U$ of $\Lambda$ is a common unit cell of $S,Q$, i.e. $S=\Lambda+(S\cap U)$ and $Q=\Lambda+(Q\cap U)$, so $\mathrm{PPC}(S)=\mathrm{PPC}(Q)$. Then all $L_{\infty}$ distances between rows in $\mathrm{PDA}(S;k),\mathrm{PDA}(Q;k)$ are the same as between the corresponding rows in $\mathrm{PDD}(S;k),\mathrm{PDD}(Q;k)$, see Definition 3.7. Hence $\mathrm{EMD}_{q}(\mathrm{PDA}(S;k),\mathrm{PDA}(Q;k))=\mathrm{EMD}_{q}(\mathrm{PDD}(S;k),\mathrm{PDD}(Q;k))\leq 2\varepsilon\sqrt[q]{k}$ by (a).
The remaining inequality follows from the $\mathrm{PDA}$ case. Indeed, each element of $\mathrm{PND}(S;k)$ in a row $i$ and a column $j=1,\dots,k$ is obtained from the corresponding element of $\mathrm{PDA}(S;k)$ by dividing by $\mathrm{PPC}(S)\sqrt[l]{j}\geq\mathrm{PPC}(S)$. Then each distance $L_{q}$ between corresponding rows in $\mathrm{PND}(S;k)$, $\mathrm{PND}(Q;k)$ is at least $\mathrm{PPC}(S)$ times smaller than between the same rows in $\mathrm{PDA}(S;k)$, $\mathrm{PDA}(Q;k)$. Then
$\mathrm{EMD}_{q}(\mathrm{PND}(S;k),\mathrm{PND}(Q;k))\leq\dfrac{\mathrm{EMD}_{q}(\mathrm{PDA}(S;k),\mathrm{PDA}(Q;k))}{\mathrm{PPC}(S)}\leq\dfrac{2\varepsilon\sqrt[q]{k}}{\mathrm{PPC}(S)}.$
Proof of Theorem 4.4.
Considering $\mathrm{PDD}(S;k)$ as a weighted distribution of rows, $\mathrm{AMD}(S;k)$ is its centroid from [cohen1997earth, section 3]. The argument below follows the proof for $q=+\infty$ of [cohen1997earth, Theorem 1] and similarly works for other invariants in parts (b,c). In the notations of Definition 4.1, we use the inequality $||\vec{u}||_{q}+||\vec{v}||_{q}||\geq||\vec{u}+\vec{v}||_{q}$ for the $q$-norm $\vec{v}||_{q}=\big(\sum\limits_{i=1}|v_{i}|^{q}\big)^{1/q}$ of the Minkowski metric $L_{q}$ as follows:
$\mathrm{EMD}_{q}(\mathrm{PDD}(S;k),\mathrm{PDD}(Q;k))=\sum\limits_{i=1}^{m(S)}\sum\limits_{j=1}^{m(Q)}f_{ij}L_{q}(\vec{R}_{i}(S),\vec{R}_{j}(Q))=$
[…]
Many authors considered criteria or complete invariants of congruence for plane quadrilaterals [vance1982minimum], whose vertices are ordered. A complete and continuous invariant of $m$ ordered points under isometry in $\mathbb{R}^{n}$ is the $m\times m$ matrix of pairwise distance [schoenberg1935remarks] or the Gram matrix of scalar products [weyl1946classical]. The extension of this approach to $m$ unordered points leads to the exponential complexity because of $m!$ permutations. For $m=4$ unordered points, Theorem 5.3 proves the completeness of $\mathrm{PDD}(C;m-1)$ under isometry in any $\mathbb{R}^{n}$. For any $m$, the invariant $\mathrm{PDD}(C;m-1)$ can computed in quadratic time $O(m^{2})$. For $m=4$, $\mathrm{PDD}(C;3)$ contains only 12 numbers (6 pairs of distances between 4 points), while $4!=24$ distance matrices on 4 points contain at least 144 numbers if we take only distances above the diagonals.
If a cloud $C$ of $m$ points has a line or plane of symmetry $L$ in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$, then all points $C\setminus L$ split into pairs of points that are symmetric in $L$ and hence have equal rows in $\mathrm{PDD}(C;m-1)$. Lemma C.5 shows that the converse holds for $m=4$.
Lemma C.5 ($\mathrm{PDD}$ detects symmetry of $m=4$ points).
For any cloud $C\subset\mathbb{R}^{n}$ of $m=4$ distinct points for $n=2,3$, if $\mathrm{PDD}(C;3)$ has two equal rows, then $C$ is either (1) mirror symmetric in the plane passing through two points of $C$ orthogonally to the line segment joining the other points of $C$, or (2) symmetric by the $180^{\circ}$ degree rotation around the line through the mid-points of two pairs of points of $C$. If $n=2$, then $C$ defines a kite, or a parallelogram or an isosceles trapezoid; see Fig. 15.
Figure 15: Left: in $\mathbb{R}^{2}$, the convex and non-convex kites have two equal rows $\{a,b,c\}$ in $\mathrm{PDD}(C;3)$ and are distinguished by $d=|p_{3}-p_{4}|$, see Lemma C.5. Middle: an isosceles trapezoid and parallelogram have $\mathrm{PDD}(C;3)$ with two pairs of equal rows $\{a,b,c\}$ and $\{a,b,d\}$, e.g. a rectangle for $c=d$. Top right: a trisosceles cloud. Bottom right: a 3-chain-equal cloud, see Example C.6.
Proof.
Let points $p_{1},p_{2}\in C$ have the same row $a\leq b\leq c$ in $\mathrm{PDD}(C;3)$. One of the distances $a,b,c$ is between the points $p_{1},p_{2}$. Without loss of generality, assume that $|p_{1}-p_{2}|=c$. Then $p_{1},p_{2}$ have distances $a,b$ to the points $p_{3},p_{4}\in C\setminus\{p_{1},p_{2}\}$.
Isosceles case. Let $|p_{1}-p_{3}|=a=|p_{2}-p_{3}|$ and $|p_{1}-p_{4}|=b=|p_{2}-p_{4}|$, see Fig. 15 (left). Then $C$ has two equal triangles $\triangle p_{1}p_{3}p_{4}=\triangle p_{2}p_{3}p_{4}$ and two isosceles triangles $\triangle p_{3}p_{1}p_{2}$ and $\triangle p_{4}p_{1}p_{2}$ with equal sides at $p_{3},p_{4}$, respectively. Let $L$ be the plane that passes through $p_{3},p_{4}$ and is orthogonal to the line segment $p_{1},p_{2}]$. Then the mirror reflection in $L$ swaps $p_{1},p_{2}$. If $n=2$, $C$ defines a (non-)convex kite.
Non-isosceles case. Then $|p_{1}-p_{3}|=a=|p_{2}-p_{4}|$ and $|p_{2}-p_{3}|=b=|p_{1}-p_{4}|$, see Fig. 15 (middle). Let $L$ be the perpendicular bisector of the line segment $p_{3},p_{4}]$. The mirror reflection in $L$ swaps $p_{3}\leftrightarrow p_{4}$ and either swaps $p_{1}\leftrightarrow p_{2}$ (then $C$ defines an isosceles trapezoid in $\mathbb{R}^{2}$) or maps $p_{2}$ to $p^{\prime}_{2}$, so that $p_{1},p^{\prime}_{2},p_{3},p_{4}$ satisfy the previous case. In the latter case, the composition with the reflection in the plane through $p_{3},p_{4}$ orthogonal to $p_{1},p^{\prime}_{2}]$ is the $180^{\circ}$ degree rotation that swaps the points as $p_{1}\leftrightarrow p_{2}$ and $p_{3}\leftrightarrow p_{4}$. If $n=2$, then $C$ defines a parallelogram, see Fig. 15 (bottom middle).
Example C.6 (trisosceles and 3-chain-equal clouds in $\mathbb{R}^{3}$).
Fig. 15 (right) shows trisosceles and 3-chain-equal clouds that have 3 pairs of equal distances and a chain of 3 equal distances, their $\mathrm{PDD}$s are $\left(\begin{array}[]{ccc}a&a&c\\ a&b&b\\ a&b&c\\ b&c&c\end{array}\right)$, $\left(\begin{array}[]{ccc}a&a&b\\ a&a&c\\ a&b&d\\ a&c&d\end{array}\right)$, respectively.
Proof of Theorem 5.3.
Case $m=2$. Any cloud $C\subset\mathbb{R}^{n}$ of $m=2$ unordered points $p_{1},p_{2}$ (labelled only for convenience) has $\mathrm{PDD}(C;1)$ consisting of the single distance $|p_{1}-p_{2}|$, which uniquely determines $C$ under isometry in any $\mathbb{R}^{n}$.
Case $m=3$. Any cloud $C\subset\mathbb{R}^{n}$ of $m=3$ unordered points with pairwise distances $a\leq b\leq c$ has $\mathrm{PDD}(C;2)=\left(\begin{array}[]{cc}a&b\\ a&c\\ b&c\end{array}\right)$. The (lexicographically) first row of $\mathrm{PDD}(C;2)$ gives us $a\leq b$. Each of the remaining two rows of $\mathrm{PDD}(C;2)$ should contain at least one value of $a$ or $b$, also in all degenerate cases such as $a=b$. Removing these repeated values from the other two rows gives us $c$, also in the case $b=c$. So $\mathrm{PDD}(C;2)$ identifies $a\leq b\leq c$ and hence $C$, uniquely under isometry in any $\mathbb{R}^{n}$.
Case $m=4$, then $n\leq 3$. For a cloud $C\subset\mathbb{R}^{3}$ of $m=4$ unordered points, $\mathrm{PDD}(C;3)$ is a $4\times 3$ matrix. Assume that $\mathrm{PDD}(C;3)$ has two equal rows $a\leq b\leq c$.
Isosceles case. In the first case of Lemma C.5 in Fig. 15 (left), $\mathrm{PDD}(C;3)$ has two more rows $\{a,a,d\}$ and $\{b,b,d\}$ including two repeated distances (say, $a,b$) among $a,b,c$. We can form two isosceles triangles with sides $a,a,c$ and $b,b,c$, which can be rotated in $\mathbb{R}^{3}$ around their common side of the length $c$, but their positions are fixed under isometry in $\mathbb{R}^{3}$ by the distance $d$ between their non-shared vertices.
Non-isosceles case. In the second case of Lemma C.5 in Fig. 15 (middle), the matrix $\mathrm{PDD}(C;3)$ has two pairs of equal rows of distances $\{a,b,c\}$ and $\{a,b,d\}$. Each of these triples uniquely determines a pair of equal triangles with a common side that are symmetric in the perpendicular bisector to this side. For example, if we start with a fixed position of $p_{3},p_{4}]$ in $\mathbb{R}^{3}$, the union of equal triangles $\triangle p_{1}p_{3}p_{4}=\triangle p_{2}p_{3}p_{4}$ in Fig. 15 (middle) is uniquely determined under isometry by the length $d$ of $p_{1},p_{2}]$. In $\mathbb{R}^{2}$, the parallelogram and isosceles trapezoid are distinguished by $d$.
Now we can assume that all rows of $\mathrm{PDD}(C;3)$ are different. Then all points can be uniquely labelled as $p_{1},p_{2},p_{3},p_{4}$ according to the lexicographic order of rows. Our aim is to get $\mathrm{PDD}(\{p_{2},p_{3},p_{4}\};2)$, reconstruct $\triangle p_{2}p_{3}p_{4}$, and then uniquely add $p_{1}$.
Case of a row with 3 equal distances. Let $\mathrm{PDD}(C;3)$ have a row of (say) $p_{1}$ with 3 equal distances $a$. After removing the row of $p_{1}$, the distance $a$ from the rows of $p_{2},p_{3},p_{4}$, we get $\mathrm{PDD}(\{p_{2},p_{3},p_{4}\};2)$. This smaller $3\times 2$ matrix determines $\triangle p_{2}p_{3}p_{4}$, uniquely under isometry in $\mathbb{R}^{3}$. For a fixed $\triangle p_{2}p_{3}p_{4}$, the position of $p_{1}$ in $\mathbb{R}^{3}$ is determined by its distance $a$ to $p_{2},p_{3},p_{4}$, uniquely under the mirror reflection relative to the plane of $\triangle p_{2}p_{3}p_{4}$. If $n=2$, then $p_{1}$ is the unique circumcenter of $\triangle p_{2}p_{3}p_{4}$.
Case of a row with 3 unique distances. Let $\mathrm{PDD}(C;3)$ have a row of (say) $p_{1}$, where each of the distances $a,b,c$ (say, to $p_{2}$, $p_{3}$, $p_{4}$) appears in at most one other row (then $a,b,c$ are distinct). After removing the row of $p_{1}$, the distance $a$ from the row $p_{2}$, the distance $b$ from the row of $p_{3}$, and the distance $c$ from the row of $p_{4}$, we get $\mathrm{PDD}(\{p_{2},p_{3},p_{4}\};2)$. This $3\times 2$ matrix determines $\triangle p_{2}p_{3}p_{4}$, uniquely under isometry in $\mathbb{R}^{3}$. Then the position of $p_{1}$ in $\mathbb{R}^{3}$ is determined by its distances $a,b,c$ to $p_{2},p_{3},p_{4}$, respectively, under a mirror reflection relative to the plane of the triangle $\triangle p_{2}p_{3}p_{4}$.
Case of one distance in 4 rows. Then two pairs of points have disjoint edges of the same length, e.g. $|p_{1}-p_{2}|=a=|p_{3}-p_{4}|$, so $\mathrm{PDD}(C;3)=\left(\begin{array}[]{ccc}a&b&c\\ a&d&e\\ a&b&d\\ a&c&e\end{array}\right)$ for $b=|p_{1}-p_{3}|$, $c=|p_{1}-p_{4}|$, $d=|p_{2}-p_{3}|$, $e=|p_{2}-p_{4}|$. Then $c\neq d$ and $b\neq e$, else $\mathrm{PDD}(C;3)$ has two equal rows (considered above), similarly when $b=c$ and $d=e$. If $a$ equals one of $b,c,d,e$ (say, $e$), $C$ is a 3-chain-equal cloud in Fig. 15 (bottom right) and the argument below still works. If $b\neq c$, we remove the row of $p_{1}$, the distance $b$ from the only row of $p_{3}$ containing $b$, the distance $c$ from the only row of $p_{4}$ containing $c$, and then remove $a$ from the remaining row of $p_{2}$. This reduction to $\mathrm{PDD}(\{p_{2},p_{3},p_{4}\};2)$ allows us to reconstruct $C$, uniquely under isometry in $\mathbb{R}^{3}$, as in the case of a row with 3 unique distances. If $b=c$ but $d\neq e$, we remove the row of $p_{2}$, the distance $d$ from the only row of $p_{3}$ containing $d$, the distance $e$ from the only row of $p_{4}$ containing $e$, and then remove $a$ from the remaining row of $p_{1}$, which allows us to uniquely reconstruct $C$, as in the case of a row with 3 unique distances above.
The final case: no distance appears in all 4 distinct rows and but every row has a distance appearing in 3 rows, hence at least four times, including two times in the same row. Then $C$ is a trisosceles cloud in Fig. 15 (top right). If any of the remaining distances $a,b,c$ are equal, $\mathrm{PDD}(C;3)$ has two equal rows (the case considered above). Then we remove any row (say $a,b,b$) with two repeated distances, the distance $b$ from the only two rows containing $b$, and the distance $a$ from the remaining row. This reduction to $\mathrm{PDD}(\{p_{2},p_{3},p_{4}\};2)$, allows us to reconstruct $\triangle p_{2}p_{3}p_{4}$, uniquely under isometry in $\mathbb{R}^{3}$. Though $p_{1}$ has equal distances to two of the vertices (say $p_{2},p_{3}$), the ambiguity of reconstructing $p_{1}$ in $\mathbb{R}^{3}$ by its distances to $p_{2},p_{3},p_{4}$, is only under the mirror reflections relative to the bisector plane of $p_{2},p_{3}]$ and the plane of $\triangle p_{2}p_{3}p_{4}$.